### v1

# Acceleration Noise Induced Decoherence in Stern-Gerlach Interferometers for Gravity Experiments

Abstract
Stern-Gerlach interferometer (SGI) is a kind of matter-wave interferometer driven by magnetic field and has been proposed for many gravity experiments. Acceleration noises such as vibration and inertial forces, together with higher-order noises like the fluctuation of the gravity gradient or the magnetic field, can cause decoherence problems of SGI, including dephasing, loss of contrast and position localization decoherence. In this paper, I will theoretically study these mechanisms of decoherence based on the analytical time-evolution operator of an SGI modeled as a harmonic oscillator under acceleration noises described by Gaussian stochastic processes.
As will be proved, for a single arm of an SGI, the shape of the Wigner function keeps invariant under an acceleration noise, although the phase and the coordinate in the classical phase space fluctuate as linear responses to the noise, and its density matrix also experiences a position localization decoherence due to the ensemble average. For the witness constructed in the spin space, the degrees of freedom in the classical phase space have to be traced out. Then acceleration noises can lead to dephasing effects on the witness, while the fluctuation on the trajectories in the phase space and position localization decoherence of the two arms will be canceled with each other. On the other hand, higher-order noises can be treated as non-uniform acceleration noises in the leading order, and they will cause decoherence in the spin space due to the contrast loss besides the dephasing. Remarkably, the randomness of the noise is essential for dephasing and position-localization decoherence, and these two mechanisms don’t cause purity loss or entropy increase if the noise is monitored as a deterministic process during the experiment.
1 Introduction
Stern-Gerlach interferometer (SGI) is a kind of matter-wave interferometer driven by an external magnetic field through the spin of the system [1]. In recent years, SGI based on Nitrogen-Vacancy (NV) centers [2] has been proposed as gravity sensors [3] like gravimeters [4, 5] and gyroscopes [6, 7], and it is also proposed for exploring the nature of gravity [8]. Compared to the matter-wave interferometers driven by laser, the test mass of an SGI can be much heavier, then the amplitude of the signal of gravitational interactions is much larger and easier to detect.
The robustness is an essential issue for all types of matter-wave interferometers including the SGI. In a real experiment, there are many sources of classical acceleration noises arising from mechanical vibrations, seismic motions, inertial forces and so on. There are also noises coupling to the test mass in different ways such as gravity gradient (tidal force) noise, rotation noises and magnetic field noise. These noises disturb the trajectories and velocities of the two arms of the interferometer at the classical level and cause decoherence in the quantum regime. There are several typical mechanisms of decoherence including the dephasing, the position localization decoherence [9, 10, 11, 12, 13, 14] and the loss of contrast (known as the Humpty-Dumpty problem [15, 16, 17]).
The dephasing effect arises from the ensemble average of the random phase fluctuations induced by the stochastic noises and results in the decoherence of the system. The dephasing can be caused by the phase fluctuation of both the path integral phase [18, 19, 20, 21] and spin phase [22]. The dephasing in the spin space has already been well-studied and is described by a decoherence rate $T_{2}^{*}$ according to the Bloch-Redfield master equation [23, 24] which of NV center is measured $T_{2}^{*}=1.68\,\rm\mu s$ [25] in experiment. On the other hand, the dephasing of the path integral phase still needs more research.
The position-localization decoherence was originally studied to understand the collapse model [9] of quantum states caused by the interactions with the ambient environment. This mechanism of decoherence is usually considered from scattering by particles like air molecules or photons [9, 26, 10], and it can be also induced by the ensemble average of stochastic noises [27, 28, 12, 14].
Both the dephasing effect and the position-localization decoherence happen in all kinds of interferometers, but the Humpty-Dumpty problem is a peculiar issue in SGI. Since the witness of an SGI is usually constructed in the spin space, the spatial Hilbert space has to be traced out, then the contrast loss of the two trajectories can lead to a decay factor of off-diagonal terms in the density matrix of the spin space, known as the Humpty-Dumpty problem. In addition, there are two typical channels of the Humpty-Dumpty problem. One is the center of mass of the two paths cannot meet with each other at the final time, and the other is the shape of the wavefunction changes.
From the math side, noise analysis for SGI is very tricky due to time dependence and randomness, so previous works usually work with some assumptions or use perturbation methods to study these effects separately. However, the Lewis-Resenfeld invariant [29, 30, 31, 32] and quantum Arnold transformation [33] provide a systematical method to solve a generic time-dependent Schrodinger equation, especially since there is an analytical solution for an SGI modeled as a harmonic oscillator under a driven force with a fluctuation. Therefore, this analytic method and the solution provide an opportunity to theoretically study all three mechanisms of decoherence in SGI at the same time.
This paper is organized as follows. In section 2, the ideal Stern-Gerlach interferometer without noises will be discussed, whereas sub-section 2.1 will study the motion of the spatial part and sub-section 2.2 will discuss the witness constructed in the spin space. The maths details for sub-sections 2.1 and 2.2 are summarized in Appendix A and B in respect. In section 3, the influence of acceleration noises on a single arm will be studied, including the dephasing, the fluctuation of the trajectory in the phase space and the position localization decoherence, which will be discussed separately in sub-sections, and the master equation will be derived in Appendix C. In section 4, the decoherence of the interferometer induced by noises will be studied, whereas sub-section 4.1 will prove that acceleration noises only have the dephasing effect, and sub-section 4.2 will treat higher-order noises as non-uniform acceleration noises and study the decoherence effects including all three mechanisms mentioned above. Finally, section 5.1 will make a discussion on the role of randomness of the noise, and section 5.2 will make a summary of this paper.
2 Ideal Stern-Gerlach interferometer without Noises
Generally speaking, the Hamiltonian of a Stern-Gerlach interferometer for gravity experiment consists of the kinetic energy $\hat{p}^{2}/2m$, the spin-magnetic field coupling $\hat{\mathbf{S}}\cdot\mathbf{B}$, the coupling with gravity and usually a trapped harmonic potential $m\omega_{0}^{2}\hat{x}^{2}/2$. For example, the 1-dimensional Hamiltonian of the gravimeter based on NV center can be written as [34]
$\hat{H}=\frac{\hat{p}^{2}}{2m}+g_{NV}\mu_{B}\frac{\partial B}{\partial x}\hat{S}_{z}\hat{x}-\frac{\chi V}{2\mu_{0}}\left(\frac{\partial B}{\partial x}\right)^{2}\hat{x}^{2}-mg\cos\theta\hat{x}-m\delta a_{n}(t)\hat{x}+\hbar D\hat{S}_{z}^{2}.$ (1)
The term $g_{NV}\mu_{B}(\partial B/\partial x)\hat{S}_{z}\hat{x}$ is the spin-magnetic field interaction, with $\mu_{B}$ is the Bohr magneton and $g_{NV}$ is the Lande g-factor of NV center. The term $\chi V/(2\mu_{0})(\partial B/\partial x)^{2}\hat{x}^{2}$ is the diamagnetic energy of the diamond, with $\chi=-2.2\times 10^{-5}$ and $V$ are the magnetic susceptibility and the volume of the diamond. The terms $mg\cos\theta\hat{x}$ and $m\delta a_{n}(t)\hat{x}$ are, respectively, the gravitational acceleration and the vibration noise. The last term $\hbar D\hat{S}_{z}^{2}$ is the zero-field splitting of the NV center with $D=(2\pi)2.8$ GHz. If the sensor is used for other gravity experiments, then the term $mg\cos\theta\hat{x}$ will change accordingly.
Since this work aims to study the spatial degrees of freedom, the spin operator $\hat{S}_{z}$ can be treated as a c-number rather than an operator for ease of calculation. Then $s_{z}=\pm 1$ of the interferometer corresponds to two paths, and each path of the interferometer is actually a trajectory of a harmonic oscillator driven by external forces. The Hamiltonian for a single arm can be written as
$\hat{H}(t)=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}-ma(t)\hat{x},$ (2)
where $ma(t)$ consists of the driven force $g_{NV}\mu_{B}(\partial B/\partial x)s_{z}\hat{x}$, the signal force $mg\cos\theta\hat{x}$ and the noise force $m\delta a_{n}(t)\hat{x}$, where the time-dependence of the Hamiltonian mainly comes from the time-dependent noise. The corresponding Schrodinger equation can be analytically solved by Lewis-Riesenfeld invariant and quantum Arnold transformation [32, 33], and the math details are summarized in A.
2.1 Time Evolution of Single Path
As is derived in A, the time-evolution operator for the Hamiltonian (2) is
$\hat{U}(t)=\mathrm{e}^{\frac{i}{\hbar}\int_{0}^{t}L[x(t^{\prime}),\dot{x}(t^{\prime})]\mathrm{d}t^{\prime}}\hat{D}(\alpha_{c}(t))\mathrm{e}^{-\frac{i}{\hbar}\hat{H}_{0}t}\hat{D}(-\alpha_{c}(0)),$ (3)
where $L[x(t),\dot{x}(t)]=\frac{1}{2}m\dot{x}^{2}-\frac{1}{2}m\omega_{0}^{2}x+ma(t)x$ is the Lagrangian corresponding to the Hamiltonian (2), $\hat{H}_{0}=\hat{p}^{2}/2m+m\omega_{0}^{2}\hat{x}^{2}/2$ is the Hamiltonian of a simple harmonic oscillator without driven forces, and $\hat{D}(\alpha_{c}(t))=\mathrm{e}^{\alpha_{c}(t)\hat{a}^{\dagger}-\alpha_{c}^{*}(t)\hat{a}}$ is the displacement operator defined by the complex coordinate $\alpha_{c}(t)=\sqrt{m\omega_{0}/2\hbar}(x_{c}(t)+ip_{c}(t)/m\omega_{0})$ of the instantaneous equilibrium position in the phase space satisfying the canonical equation in complex coordinate
$\dot{\alpha}_{c}=i\omega_{0}\alpha_{c}+i\sqrt{\frac{m}{2\hbar\omega_{0}}}a(t).$ (4)
The solution of $\alpha_{c}(t)$ is
$\alpha_{c}(t)=i\sqrt{\frac{m}{2\hbar\omega_{0}}}\mathrm{e}^{i\omega_{0}t}\int_{0}^{t}a(t^{\prime})\mathrm{e}^{-i\omega_{0}t^{\prime}}\mathrm{d}t^{\prime}.$ (5)
The physical interpretation of the displacement operator $\hat{D}(\alpha_{c}(t))$ is a reference transformation in the phase space from the lab reference frame to the co-moving reference of the equilibrium position of the oscillator. Thus, the physical intuition of the time evolution operator $\hat{U}(t)$ given by (3) is that the system is a simple oscillator without any driven force in the co-moving reference frame of the equilibrium position. Furthermore, there is a Feynman’s path integral phase $\mathrm{e}^{\frac{i}{\hbar}\int_{0}^{t}L[x(t^{\prime}),\dot{x}(t^{\prime})]\mathrm{d}t^{\prime}}$ along the classical trajectory of the equilibrium position of the oscillator.
For an arbitrary initial state described by a density matrix $\hat{\rho}_{0}$, it evolves as $\hat{\rho}(t)=\hat{U}(t)\hat{\rho}_{0}\hat{U}^{\dagger}(t)$. From the point of view of Weyl quantization, the Wigner function is the Weyl-Wigner transform of the density matrix [35], i.e. $W(\alpha)=\int\text{Tr}[\hat{\rho}\hat{D}(\alpha^{\prime})]\mathrm{e}^{\alpha^{\prime*}\alpha-\alpha^{\prime}\alpha^{*}}/\pi^{2}\mathrm{d}^{2}\alpha^{\prime}$. Then as is derived in A, the time-evolution of the Wigner function of an arbitrary state satisfies
$W(\alpha,t)=W(\alpha_{c}(0)+\mathrm{e}^{i\omega_{0}t}(\alpha-\alpha_{c}(t)),t=0).$ (6)
Thus, the shape of the Wigner function of an arbitrary state doesn’t change during time evolution, and its time evolution is only a transformation in the phase space composited by a translation $\alpha_{c}(t)$ plus a rotation with a constant angular speed $\omega_{0}$. Besides, the evolution of an arbitrary quantum state is totally determined by its classical trajectory $x_{c}(t)$ and momentum $p_{c}(t)$ satisfying the classical equations of motion (4) in the phase space.
In an experiment, the initial state is usually prepared in a coherent state |$\ket{\beta_{0}}=\hat{D}(\beta_{0})\ket{0}$, then the time-evolution (3) for the coherent state |$\ket{\psi(0)}=\ket{\beta_{0}}$ is  [^1] According to the Baker-Campbell-Hausdorff (BCH) formula, one may check an identity for the displacement operator as $\hat{D}(\alpha+\beta)=\hat{D}(\alpha)\hat{D}(\beta)\mathrm{e}^{-(\alpha\beta^{*}-\alpha^{*}\beta)/2}$. Then one may obtain a property for the displacement operator acting on a coherent state as $\hat{D}(\alpha)\ket{\beta}=\mathrm{e}^{-(\alpha\beta^{*}-\alpha^{*}\beta)/2}\ket{\alpha+\beta}=\mathrm{e}^{-i\mathrm{Im}(\alpha\beta^{*})}\ket{\alpha+\beta}$, where there is an identity $\alpha\beta^{*}-\alpha\beta^{*}=2i\mathrm{Im}(\alpha\beta^{*})$. Note that the time-evolution of the coherent state is $\mathrm{e}^{-i\frac{\hat{H}t}{\hbar}}\ket{\alpha+\beta}=\mathrm{e}^{-\frac{1}{2}i\omega_{0}t}\ket{\mathrm{e}^{-i\omega_{0}t}(\alpha+\beta)}$. Combining these properties of displacement operators, one can compute the time evolution of a coherent state and obtain (7).
|$\ket{\psi(t)}=\mathrm{e}^{\frac{i}{\hbar}\int_{0}^{t}L[x_{c}(t^{\prime}),\dot{x}_{c}(t^{\prime})]\mathrm{d}t^{\prime}}\mathrm{e}^{-i\mathrm{Im}(\Delta\alpha_{c}\beta_{0}^{*})}\mathrm{e}^{-\frac{1}{2}\omega_{0}t}\ket{\alpha_{c}(t)+\mathrm{e}^{-i\omega_{0}t}(\beta_{0}-\alpha_{c}(0))},$ (7)
where $\Delta\alpha_{c}\equiv\alpha_{c}(t)-\alpha_{c}(0)$. As is shown in (7), the evolution of a single arm is always a coherent state, so the shape of the wavefunction keeps a Gaussian wave-packet with a fixed width $\sqrt{2\hbar/m\omega_{0}}$ during evolution. Furthermore, the total phase consists of a Feynman’s path integral phase (or a classical phase) $\int_{0}^{t}L[x_{c}(t^{\prime}),\dot{x}_{c}(t^{\prime})]/\hbar\mathrm{d}t^{\prime}$ and additional quantum phases $\mathrm{Im}(\Delta\alpha_{c}\beta_{0}^{*})$ and $-\omega_{0}t/2$ related to the details of the initial quantum state.
As an example, consider a constant driven force $F(t)\equiv ma_{0}$ without noises and an initial state prepared in a coherent state |$\ket{\psi(0)}=\ket{\beta_{0}}$ with $x_{c}(0)=0$ and $p_{c}(0)=0$, the instantaneous equilibrium position satisfies $x_{c}(t)\equiv a_{0}/\omega_{0}^{2}$ and $\dot{x}_{c}(t)\equiv 0$ for $t>0$, then the time-evolution of the state is
|$\ket{\psi(t)}=\mathrm{e}^{\frac{i}{2\hbar}m\frac{a_{0}^{2}}{\omega_{0}^{2}}t}\mathrm{e}^{-\frac{1}{2}\omega_{0}t}\Ket{\sqrt{\frac{m\omega_{0}}{2\hbar}}\frac{a_{0}}{\omega_{0}^{2}}+\left(\beta_{0}-\sqrt{\frac{m\omega_{0}}{2\hbar}}\frac{a_{0}}{\omega_{0}^{2}}\right)\mathrm{e}^{-i\omega_{0}t}}.$ (8)
This result is the same as the result in [4, 36] except a phase proportional to $\sin\omega_{0}t$ which is related to the spin-orbit coupling $\hat{S}_{z}\hat{x}$. It is because the spin is treated as a c-number rather than an operator in (1) and (2). But this phase is independent of spatial degrees of freedom, so acceleration noises don’t affect it. Hence, this phase will be neglected in the paper.
2.2 Witness of Stern-Gerlach Interferometer
Based on the wavefunction of a single arm, a Stern-Gerlach interferometer is usually initially prepared as |$\ket{\Psi(0)}=\frac{1}{\sqrt{2}}(\ket{S_{z}=1}+\ket{S_{z}=-1})\ket{\beta_{0}}$, and the states with $S_{z}=\pm 1$ experience different driven forces in (2) according to (1). Then the corresponding spatial parts evolve along different trajectories described by their complex coordinates $\beta_{1}(t)$ and $\beta_{-1}(t)$ in the classical phase space according to (7), that is
$\beta_{s_{z}}(t)=\alpha_{s_{z}}(t)+\mathrm{e}^{-i\omega_{0}t}\left(\beta_{0}-\alpha_{s_{z}}(0)\right),$ (9)
with $s_{z}=\pm 1$. Furthermore, the integral phases $\phi_{1}(t)$ and $\phi_{-1}(t)$ of the two paths also evolve differently according to (7). By designing the magnetic field, the spatial parts for $S_{z}=\pm 1$ can meet with each other at the final time $t_{f}$, i.e. $\beta_{1}(t_{f})=\beta_{-1}(t_{f})\equiv\beta(t_{f})$. Then the density matrix $\hat{\rho}=\ket{\Psi(t_{f})}\bra{\Psi(t_{f})}$ | of the final state is
[…]
where the differential phase is defined as $\phi_{\rm diff}\equiv\phi_{-1}(t_{f})-\phi_{1}(t_{f})$. Since the two trajectories meet with each other at $t_{f}$, the quantum phases related to $\alpha_{c}(t_{f})$ in $\phi_{1}(t_{f})$ and $\phi_{-1}(t_{f})$ will cancel with each other, so the differential phase $\phi_{\rm diff}$ only relies on the path integral phase.
The final observable quantity of the Stern-Gerlach interferometer is usually constructed by a witness operator $\hat{W}$ in the spin space, so the spacial wavefunctions have to be traced out, i.e. $\hat{\rho}_{\rm spin}(t)=\mathrm{Tr}_{\beta}[\hat{\rho}(t)]$. Then the density matrix in spin space at the final time is simply
$\hat{\rho}_{\rm spin}(t_{f})=\frac{1}{2}\begin{pmatrix}1&0&\mathrm{e}^{-i\phi_{\rm diff}}\\ 0&0&0\\ \mathrm{e}^{i\phi_{\rm diff}}&0&1\end{pmatrix}.$ (11)
A common-used witness to measure the differential phase $\phi_{\rm diff}$ is the Ramsey interferometry [37, 4, 23]. It is constructed by measuring the population of |$\ket{S_{z}=0}$ state after a $\pi/2$-pulse operating on the state. As is derived in the B, the witness operator for Ramsey interferometry is given by
$\hat{W}=\begin{pmatrix}\frac{1}{2}&0&\frac{1}{2}\\ 0&0&0\\ \frac{1}{2}&0&\frac{1}{2}\end{pmatrix}.$ (12)
Then one can obtain the observable result for the Ramsey interferometry is $\mathrm{Tr}[\hat{\rho}_{\rm spin}(t_{f})\hat{W}]=\cos^{2}(\phi_{\rm diff}/2)$, which means that the population of the eigenstate |$\ket{S_{z}=0}$ at the final time after a $\pi/2$-pulse is $\cos^{2}(\phi_{\rm diff}/2)$. For a Stern-Gerlach interferometer based on NV center described by the Hamiltonian (1), the differential phase $\phi_{\rm diff}$ is proportional to the gravitational acceleration $g$, so this device can be used as a gravimeter to measure the local gravity $g$ through the witness (12) of Ramsey interferometry [4].
3 Single Arm under Acceleration Noises
In a real experiment, there always exist some acceleration noises $\delta a(t)$ such as the vibration of the mechanical support of the system, the inertial force due to the vibration of the lab (known as the residue acceleration noise (RAN) [18, 19]) and electric forces caused by ambient electric particles [20].
There are also some other noises coupling to the interferometer in different ways such as the gravity gradient noise (GGN) described by $m\delta\omega^{2}(t)\hat{x}^{2}/2$ [18, 19] and the fluctuation of the magnetic field described by $g\mu_{B}\delta\eta(t)\hat{S}_{z}\hat{x}$ with $\eta$ denoting the gradient of the magnetic field $\eta\equiv\partial B(t)/\partial x$. But these noises can be approximately regarded as acceleration noises by treating some operators as c-numbers. In particular, one may define $\delta a(x,t)=\delta\omega^{2}(t)x/2$ and $\delta a(s_{z},t)=g\mu_{B}s_{z}\delta\eta(t)/m$ with a c-number $s_{z}=\pm 1$, then the Hamiltonians for GGN and magnetic field noise become $m\delta a(x,t)\hat{x}$ and $m\delta a(s_{x},t)\hat{x}$. Thus, the theory for acceleration noises can be used to study the leading-order effects of these noises.
In most experimental situations, acceleration noises can be described by a (wide-sense) stationary Gaussian process with a zero mean value. In particular, at any time $t$, the acceleration noise $\delta a(t)$ follows a Gaussian distribution. In addition, the wide-sense stationarity requires the auto-correlation function of $\delta a(t)$ doesn’t vary with respect to time $t$, i.e. $\mathbb{E}[\delta a(t)\delta a(t+\tau)]$ only depends on $\tau$, where $\mathbb{E}[\cdot]$ represents the statistical average of stochastic variables. According to the Wiener-Khinchin theorem, the auto-correlation function of $\delta a(t)$ is the Fourier transform of the corresponding power spectrum density (PSD)
$\mathbb{E}[\delta a(t)]$ $=0,\quad\forall t,$ (13)
$\mathbb{E}[\delta a(t)\delta a(t+\tau)]$ $=\int S_{aa}(\omega)\mathrm{e}^{-i\omega\tau}\mathrm{d}\omega,\quad\forall t,$
where $S_{aa}(\omega)=\lim\limits_{T\to\infty}|\delta a_{T}(\omega)|^{2}/T$ and $\delta a_{T}(\omega)$ is the Fourier transform of $\delta a(t)$ on the finite time domain $0\sim T$. Note that the second equation is independent with $t$ because the noise is assumed stationary. In addition, let $\tau=0$, then the auto-correlation function of $\delta a(t)$ gives the variance of $\delta a(t)$ as $\sigma_{a}^{2}(t)\equiv\mathbb{E}[(\delta a(t))^{2}]=\int S_{aa}(\omega)\mathrm{d}\omega$, which is also time-independent.
For a single arm, the acceleration noise $\delta a(t)$ is part of the $-ma(t)\hat{x}$ term in (2), so the time evolution operator (3) also holds when the noise is present, although the phase and the complex coordinate $\alpha_{c}(t)$ fluctuate. As a consequence, the displacement operators $\hat{D}(\alpha_{c}(0))$ keeps the same and $\hat{D}(\alpha_{c}(t))\to\hat{D}(\alpha_{c}(t)+\delta\alpha)$. Thus, the time evolution of the Wigner function of an arbitrary initial state becomes
$W(\alpha,t)=W(\alpha_{c}(0)+\mathrm{e}^{i\omega_{0}t}(\alpha-\alpha_{c}(t)-\delta\alpha),t=0).$ (14)
The first notable conclusion is that the acceleration noise $\delta a(t)$ doesn’t change the shape of the Wigner function of an arbitrary initial state, and it only affects the path integral phase $\delta\phi$ and the classical trajectory $\delta\alpha$ in phase space. Specifically, for a coherent state given by (7), the wavefunction under the acceleration noise is
[…]
As is proved for general cases, the wavefunction is still a coherent state under the noise $\delta a(t)$, and there are fluctuations of the phase $\delta\phi$ and the complex coordinate $\delta\alpha$ in the phase space.
3.1 Fluctuation of path integral phase $\delta\phi$
Generally speaking, acceleration noises can fluctuate the path integral phase through two channels at the leading order. The first channel is through the path integral of the Lagrangian $m\delta a(t)x$ along the classical trajectory $x_{c}(t)$, and the second is through the fluctuation of the classical trajectory $\delta x$ and velocity $\delta\dot{x}$ in the path integral of the ideal Lagrangian. The combination of these two channels is a higher-order effect and can be neglected. Then the phase fluctuation can be written as
$\delta\phi=\frac{1}{\hbar}\int m\dot{x}\delta\dot{x}-m\omega_{0}^{2}x\delta x+ma(t)\delta x+m\delta a(t)x\mathrm{d}t.$ (16)
It is notable that the dynamical equation (4) can be written as another form as $\ddot{x}+\omega_{0}^{2}x=a(t)$, then one can find out an identity
$m\dot{x}\delta\dot{x}-m\omega_{0}^{2}x\delta x+ma(t)\delta x=m\frac{\mathrm{d}}{\mathrm{d}t}\left(\dot{x}\delta x\right).$ (17)
Thus, the phase fluctuation arising from the fluctuation of the trajectory $\delta x(t)$ and velocity $\delta\dot{x}(t)$ is a total derivative that will vanish after the path integral. Remarkably, some authors have already noticed that this channel of the phase fluctuation has to be taken into account in principle, but they neglected this effect by assuming it is a higher-order effect [18, 19, 20, 21]. As is proved above, the dynamical equation ensures this effect is strictly negligible without any assumptions  [^2] This conclusion can be generalized to a generic matter-wave interferometer under a general noise. Suppose the interferometer is built up through a Lagrangian $L[x_{i}(t),\dot{x}_{i}(t)]$ and the noise is described by a Lagrangian $\delta L[x_{i}(t),\dot{x}_{i}(t)]$. Denote the ideal trajectory as $x_{i}^{0}(t)$, and the ideal phase is $\phi^{0}=\int L[x_{i}^{0},\dot{x}_{i}^{0}]/\hbar\mathrm{d}t$. Then the phase under noise at the leading order is $\begin{split}\phi^{0}+\delta\phi&=\frac{1}{\hbar}\int L[x_{i}^{0}+\delta x_{i},\dot{x}_{i}^{0}+\delta\dot{x}_{i}^{0}]+\delta L[x_{i}^{0},\dot{x}_{i}^{0}]\mathrm{d}t\\ &=\frac{1}{\hbar}\int L[x_{i}^{0},\dot{x}_{i}^{0}]+\frac{\partial L}{\partial x_{i}}\bigg{|}_{(x_{i}^{0},\dot{x}_{i}^{0})}\delta x_{i}+\frac{\partial L}{\partial\dot{x}_{i}}\bigg{|}_{(x_{i}^{0},\dot{x}_{i}^{0})}\delta\dot{x}_{i}+\delta L[x_{i}^{0},\dot{x}_{i}^{0}]\mathrm{d}t.\end{split}$ According to the Euler-Lagrangian equation $\partial L/\partial x_{i}=\mathrm{d}(\partial L/\partial\dot{x}_{i})/\mathrm{d}t$, one can prove an identity $\frac{\partial L}{\partial x_{i}}\bigg{|}_{(x_{i}^{0},\dot{x}_{i}^{0})}\delta x_{i}+\frac{\partial L}{\partial\dot{x}_{i}}\bigg{|}_{(x_{i}^{0},\dot{x}_{i}^{0})}\delta\dot{x}_{i}=\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial L}{\partial\dot{x}_{i}}\bigg{|}_{(x_{i}^{0},\dot{x}_{i}^{0})}\delta x_{i}\right).$ Thus, the phase fluctuation arising from the fluctuation of the trajectory $\delta x_{i}$ and the velocity $\delta\dot{x}_{i}$ is the integral of a total derivative. Then one may read $\delta\phi=\int\delta L[x_{i}^{0},\dot{x}_{i}^{0}]/\hbar\mathrm{d}t$ by neglecting this total derivative term. . Therefore, the phase fluctuation is simply
$\delta\phi(t)=\frac{1}{\hbar}\int_{0}^{t}m\delta a(t^{\prime})x(t^{\prime})\mathrm{d}t^{\prime}.$ (18)
So the phase fluctuation $\delta\phi(t)$ is an Ito integral of the stochastic noise $\delta a(t)$ and can be regarded as a linear response to $\delta a(t)$. Since the acceleration noise $\delta a(t)$ is assumed as a stationary Gaussian process, the phase fluctuation $\delta\phi(t)$ is a Gaussian process with a zero mean-value $\mathbb{E}[\delta\phi(t)]=0$ for arbitrary time $t$, although it is no longer stationary. The variance of $\delta\phi(t)$ can be computed as
$\sigma_{\phi}^{2}(t)\equiv\mathbb{E}[\left(\delta\phi(t)\right)^{2}]=\frac{m^{2}}{\hbar^{2}}\int S_{aa}(\omega)|x_{t}(\omega)|^{2}\mathrm{d}\omega,$ (19)
where $x_{t}(\omega)$ is the Fourier transform of the trajectory $x(t^{\prime})$ in the finite time domain $0\sim t$, which can be written by a convolution between $x(\omega)$ and the sinc-function as $x_{t}(\omega)=x(\omega)\ast\text{sinc}(\omega t)$. Furthermore, $x(\omega)$ plays a role of the transfer function or frequency response function of the linear response of $\delta\phi$ to $\delta a$ [38, 19, 21].
3.2 Fluctuation of classical phase space $\delta\alpha$
As is shown in (15), the acceleration noise $\delta a(t)$ can also cause a fluctuation of the coordinate $\delta\alpha$ in phase space. According to the time-evolution (5) of the complex coordinate $\alpha_{c}$ in the phase space, its fluctuation $\delta\alpha$ under the acceleration noise $\delta a(t)$ is given by
$\delta\alpha(t)=i\sqrt{\frac{m}{2\hbar\omega_{0}}}\mathrm{e}^{i\omega_{0}t}\int_{0}^{t}\delta a(t^{\prime})\mathrm{e}^{-i\omega_{0}t^{\prime}}\mathrm{d}t^{\prime}.$ (20)
So $\delta\alpha(t)$ is also a linear response to the acceleration noise $\delta a(t)$. Besides, $\delta\alpha(t)$ is also a Gaussian process with a zero mean-value $\mathbb{E}[\delta\alpha(t)]=0$ and its auto-correlation function is
$\mathbb{E}[\delta\alpha(t)\delta\alpha^{*}(t+\tau)]=\frac{m}{2\hbar\omega_{0}}\mathrm{e}^{-\frac{i}{2}\omega_{0}\tau}t(t+\tau)\int S_{aa}(\omega)\mathrm{e}^{-\frac{i}{2}\omega\tau}\text{sinc}\frac{(\omega-\omega_{0})t}{2}\text{sinc}\frac{(\omega-\omega_{0})(t+\tau)}{2}\mathrm{d}\omega,$ (21)
Remarkably, $\mathbb{E}[\delta\alpha(t)\delta\alpha^{*}(t+\tau)]$ relies on both $t$ and $\tau$, so $\delta\alpha$ is not stationary. Let $\tau=0$, then one may have the variance of $\delta\alpha$ at any time $t$ as
$\sigma_{\alpha}^{2}(t)\equiv\mathbb{E}[|\delta\alpha(t)|^{2}]=\frac{m}{2\hbar\omega_{0}}t^{2}\int S_{aa}(\omega)\text{sinc}^{2}\frac{(\omega-\omega_{0})t}{2}\mathrm{d}\omega.$ (22)
Thus, $\delta\alpha$ is a kind of diffusion process. During the beginning of the diffusion process  [^3] Here, asymptotic behaviours of the sinc-function have been used. That is, $\text{sinc}^{2}((\omega-\omega_{0})t/2)=1+\mathcal{O}(t^{2})$ for small $t$, and $\text{sinc}^{2}((\omega-\omega_{0})t/2)=2\delta(\omega-\omega_{0})/(\pi t)$ for large $t$., $\delta\alpha$ diffuses as $\sigma_{\alpha}^{2}(t)=D_{1}t^{2}$ with a generalized diffusion parameter $D_{1}=m\sigma_{a}^{2}/(2\hbar\omega_{0})$. In the long time limit, $\delta\alpha$ diffuses as $\sigma_{\alpha}^{2}(t)=D_{2}t$, which is similar to the standard Brownian motion with a diffusion parameter $D_{2}=mS_{aa}(\omega_{0})/(\pi\hbar\omega_{0})$. In other words, the diffusion process of $\delta\alpha$ in the long time limit is Markovian or memoryless, which means its long-time diffusion is independent of the history of the beginning diffusion.
In frequency space, the dynamical equation (4) of $\alpha_{c}$ implies a frequency response relation as $\delta\alpha(\omega)=\chi(\omega)\delta a(\omega)$, where the factor $\chi(\omega)=\sqrt{m/(2\hbar\omega_{0})}/(\omega-\omega_{0})$ is a generalized mechanical susceptibility of $\alpha_{c}(\omega)$. Then the PSD of $\delta\alpha$ is
$S_{\alpha\alpha}(\omega)=\frac{m}{2\hbar\omega_{0}}\frac{S_{aa}(\omega)}{(\omega-\omega_{0})^{2}}.$ (23)
It is notable that the Wiener-Khinchin theorem doesn’t hold for $\delta\alpha$ because it is not stationary, and one may check that the relationship between the auto-correlation function and PSD of $\delta\alpha$ is
$\mathbb{E}[\delta\alpha(t)\delta\alpha^{*}(t+\tau)]=\int S_{\alpha\alpha}(\omega)\mathrm{e}^{-i\omega\tau}\left(1+\mathrm{e}^{-i(\omega-\omega_{0})\tau}-\mathrm{e}^{i(\omega-\omega_{0})t}-\mathrm{e}^{-i(\omega-\omega_{0})(t+\tau)}\right)\mathrm{d}\omega.$ (24)
3.3 Position localization decoherence
Since the trajectory $\alpha_{c}(t)$ in the classical phase space is stochastic, the wavefunction (15) is also random. This randomness will cause a position localization decoherence [10] on the single arm of the interferometer [12, 13]. In particular, the density matrix $\hat{\rho}(t)=\ket{\psi(t)}\bra{\psi(t)}$ | for the coherent state (15) under the acceleration noise is
$\bra{x_{1}}\hat{\rho}(t)\ket{x_{2}}=\sqrt{\frac{m\omega_{0}}{\pi\hbar}}\exp\left[-\frac{m\omega_{0}}{2\hbar}(x_{1}^{2}+x_{2}^{2})+\sqrt{\frac{2m\omega_{0}}{\hbar}}\mathrm{Re}\beta(t)(x_{1}+x_{2})+i\sqrt{\frac{2m\omega_{0}}{\hbar}}\mathrm{Im}\beta(t)(x_{1}-x_{2})-2(\mathrm{Re}\beta(t))^{2}\right],$ (25)
where the notation $\beta(t)\equiv\alpha_{c}(t)+\delta\alpha(t)+\mathrm{e}^{-i\omega_{0}t}(\beta-\alpha_{c}(0))$ is the complex coordinate of the coherent state under the acceleration noise $\delta a(t)$ in the classical phase space. Then the ensemble average of the density matrix satisfies
$\bra{x_{1}}\mathbb{E}[\hat{\rho}(t)]\ket{x_{2}}=\bra{x_{1}}\hat{\rho}(0)\ket{x_{2}}\mathrm{e}^{-\frac{m\omega_{0}}{\hbar}(x_{1}-x_{2})^{2}\sigma_{\text{Im}\alpha}^{2}(t)},$ (26)
where the terms related to $\mathrm{Re}\beta(t)$ vanish and $\sigma_{\text{Im}\alpha}^{2}(t)$ is the variance of $\text{Im}\beta(t)$. In addition, since the variance of $\delta\alpha$ is approximately proportional to $t$ for the long time limit, then there is a well-defined decoherence rate $\gamma=m^{2}(x_{1}-x_{2})^{2}S_{aa}(\omega_{0})/(\pi\hbar^{2})$. It is noteworthy that a decoherence rate proportional to $(x_{1}-x_{2})^{2}$ is related to the following master equation [10, 12]
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{\rho}=\frac{1}{i\hbar}[\hat{H},\hat{\rho}]-\Lambda[\hat{x},[\hat{x},\hat{\rho}]],$ (27)
where $\Lambda=m^{2}S_{aa}(\omega_{0})/(\pi\hbar^{2})$, which is the same as [12] upto a factor $2/\pi$. Such master equation can be derived directly from the Hamiltonian (2), and the math details are summarized in C. Remarkably, the master equation (27) is exactly the long-wavelength limit of the decoherence due to the environmental scattering [26, 10]. Therefore, the scattering of long-wavelength particles like photons can be regarded as a classical stochastic noise in a long-time limit.
4 Interferometer under Acceleration Noises
Figure 1: The trajectories in phase space affected by acceleration noises, where (a) for uniform noises and (b) for non-uniform noises. As is shown in (a), under uniform noises, the final points in phase space of both arms translate the same, so the two paths can be overlapped. As is shown in (b), under non-uniform noises, the points of both arms translate differently, so there is a contrast loss as $|\braket{\alpha_{1}}{\alpha_{2}}|$ |.
As discussed in the previous section, acceleration noises don’t change the shape of the wavefunction of every arm of the interferometer. However, they can fluctuate the phase and the classical trajectory of each arm as linear responses and also induce position localization decoherence on each arm.
For the interferometer with two arms, the noise can act on the two arms in the same or different ways. For the noises originating from random acceleration such as vibration or inertial force, the two arms experience the same random acceleration and the trajectories can meet with each other at the final time, shown as Fig.1 (a). For the noises arising from the fluctuation of gravity gradient or magnetic field gradient, the fluctuations on the two arms are different, so they are not able to meet with each other, shown as Fig.1 (b). As will be discussed as follows, uniform acceleration noises only cause dephasing, while non-uniform acceleration additionally causes the Humpty-Dumpty problem and position localization decoherence.
4.1 Acceleration noises
Under a uniform acceleration noise, the quantum state at the final time under acceleration noises becomes
|$\ket{\Psi(t_{f})}=\frac{1}{\sqrt{2}}(\mathrm{e}^{i(\phi_{1}(t_{f})+\delta\phi_{1})}\ket{S_{z}=1}+\mathrm{e}^{i(\phi_{-1}(t_{f})+\delta\phi_{-1})}\ket{S_{z}=-1})\ket{\beta(t_{f})+\delta\alpha}.$ (28)
Thus, the spin space and the coordinate space are still separable, so the fluctuation on |$\ket{\beta(t_{f})+\delta\alpha}$ doesn’t affect the density matrix of the spin space after tracing out. Therefore, the density matrix in spin space is
$\mathbb{E}[\hat{\rho}_{\rm spin}(t_{f})]=\frac{1}{2}\begin{pmatrix}1&0&\mathrm{e}^{-\sigma_{\phi_{\rm diff}}^{2}}\mathrm{e}^{-i\phi_{\rm diff}}\\ 0&0&0\\ \mathrm{e}^{-\sigma_{\phi_{\rm diff}}^{2}}\mathrm{e}^{i\phi_{\rm diff}}&0&1\end{pmatrix},$ (29)
where $\phi_{\rm diff}=\phi_{-1}(t_{f})-\phi_{1}(t_{f})$ is the differential phase and $\delta\phi_{\rm diff}$ is its fluctuation. As discussed in the previous section, the phase fluctuations of both arms follow Gaussian distributions, so the exponential decay factor $\mathrm{e}^{-\sigma^{2}_{\rm diff}}$ arises from the ensemble average of the random phase factor, i.e. $\mathbb{E}[\mathrm{e}^{i\delta\phi_{\rm diff}}]=\int\mathrm{e}^{i\delta\phi_{\rm diff}}p(\delta\phi_{\rm diff})\mathrm{d}\delta\phi_{\rm diff}=\mathrm{e}^{-\sigma^{2}_{\rm diff}}$, where $p(\delta\phi_{\rm diff})=\mathrm{e}^{-\delta\phi^{2}_{\rm diff}/(2\sigma^{2}_{\rm diff})}/(\sqrt{2\pi}\sigma_{\rm diff})$ is the Gaussian probablity density function for $\delta\phi_{\rm diff}$. In addition, the phase fluctuations of both arms are linear responses to the acceleration noises, so the differential phase is also a linear response to $\delta a$ and its variance is
$\delta\phi_{\rm diff}$ $=\frac{m}{\hbar}\int_{0}^{t_{f}}\delta a(t)(x_{1}(t)-x_{-1}(t))\mathrm{d}t,$ (30)
$\sigma_{\phi_{\rm diff}}^{2}$ $=\frac{m^{2}}{\hbar^{2}}\int S_{aa}(\omega)F(\omega)\mathrm{d}\omega,$
where $F(\omega)=|x_{1}(\omega)-x_{-1}(\omega)|^{2}$ is the transfer function [38, 19, 21]. Generally speaking, the transfer function $F(\omega)$ tends constant in the low-frequency limit and decreases as $\omega^{-2k}$ for some index $k$ in the high-frequency limit. In particular, in the low-frequency limit $\omega\to 0$, the factor $\mathrm{e}^{i\omega t}\to 1$, then the Fourier transform satisfies $x(\omega)=\int x(t)\mathrm{e}^{i\omega t}\mathrm{d}t\approx\int x(t)\mathrm{d}t$ which is independent with $\omega$, so $F(\omega)$ is approximately constant to $\omega$ in the low-frequency limit. In the high-frequency limit, since the trajectories can be expanded as Taylor’s series as $x_{1}(t)-x_{-1}(t)=\sum_{n=k}^{\infty}c_{n}t^{n}$ with the leading order as $t^{k}$, then the corresponding Fourier transform is $x_{1}(\omega)-x_{-1}(\omega)=\sum_{n=k}^{\infty}c_{n}\omega^{-n-1}$, so the transfer function is proportional to $\omega^{-2k-2}$ at the leading order in the high-frequency limit.
4.2 Non-uniform noises
Apart from acceleration noises, there exist noises coupling to the interferometer in other ways such as $m\delta\omega^{2}(t)\hat{x}^{2}$ or $\mu\delta\eta(t)\hat{S}_{z}\hat{x}$ with $\eta(t)\equiv\partial B(t)/\partial x$. In order to study these noises analytically, it is necessary to construct the corresponding Lewis-Riesenfeld invariant for these coupling Hamiltonians, which can be a great challenge in math. However, by treating some operators as c-numbers, these noises can be regarded as non-uniform acceleration noises. Then the effects of these noises can be studied through the analytic solution for acceleration noises, although some higher-order properties can be lost.
As is shown in Fig.1 (b), the final states of the two paths are usually not able to meet with each other, i.e. $\beta_{1}(t_{f})\neq\beta_{-1}(t_{f})$. Then the quantum state at the final time under acceleration noises becomes
|$\ket{\Psi(t_{f})}=\frac{1}{\sqrt{2}}(\mathrm{e}^{i(\phi_{1}(t_{f})+\delta\phi_{1})}\ket{S_{z}=1}\ket{\beta_{1}}+\mathrm{e}^{i(\phi_{-1}(t_{f})+\delta\phi_{-1})}\ket{S_{z}=-1}\ket{\beta_{-1}}),$ (31)
where $\beta_{s_{z}}\equiv\beta(t_{f})+\delta\alpha_{s_{z}}$ for $s_{z}=\pm 1$ are the complex coordinates in the phase space for the two arms under noises. As is discussed in Section 3, the phase fluctuations $\delta\phi_{s_{z}}$ and the complex coordinate fluctuations $\delta\alpha_{s_{z}}$ are linear responses to the acceleration noises and follow Gaussian distributions.
Since the witness is constructed in the spin space, the spatial parts in the density matrix have to be traced out. In addition, since the fluctuation of the phases $\delta\phi_{s_{z}}$ and the trajectories $\delta\alpha_{s_{z}}$ are stochastic, the density matrix also needs to be ensemble-averaged out, then the density matrix becomes
$\mathbb{E}[\hat{\rho}_{\rm spin}(t_{f})]=\frac{1}{2}\begin{pmatrix}1&0&\mathbb{E}[\braket{\beta_{-1}}{\beta_{1}}]\mathrm{e}^{-\sigma_{\phi_{\rm diff}}^{2}}\mathrm{e}^{-i\phi_{\rm diff}}\\ 0&0&0\\ \mathbb{E}[\braket{\beta_{1}}{\beta_{-1}}]\mathrm{e}^{-\sigma_{\phi_{\rm diff}}^{2}}\mathrm{e}^{i\phi_{\rm diff}}&0&1\end{pmatrix}.$ (32)
Comparing the ideal density matrix (11), the off-diagonal terms decay due to two factors $\mathrm{e}^{-\sigma_{\phi_{\rm diff}}^{2}}$ and $\mathbb{E}[\braket{\beta_{-1}}{\beta_{1}}]$ corresponding to the dephasing and Humpty-Dumpty problem in respect. Therefore, the dephasing and contrast loss in the spatial part of the interferometer induce decoherence in the spin space.
The fluctuation of the differential phase of the interferometer is given by
$\delta\phi_{\rm diff}=\frac{m}{\hbar}\int_{0}^{t_{f}}(\delta a_{1}(t)x_{1}(t)-\delta a_{-1}(t)x_{-1}(t))\mathrm{d}t,$ (33)
and its variance is determined by the correlation between $\delta a_{1}(t)$ and $\delta a_{-1}(t)$. For example, if $\delta a_{1}(t)$ and $\delta a_{-1}(t)$ arise from a noise term $\delta L=mf(t)x^{n}$ for some Gaussian noise $f(t)$ by defining $\delta a_{s_{z}}(t)=f(t)x_{s_{z}}^{n-1}(t)$ with $s_{z}=\pm 1$, then the dephasing of the single arm becomes
$\delta\phi_{\rm diff}=\frac{m}{\hbar}\int_{0}^{t_{f}}f(t)x^{n}(t)\mathrm{d}t.$ (34)
Thus, the variance of the differential phase also has a form (30) with a transfer function $F(\omega)=|x_{1}^{n}(\omega)-x_{-1}^{n}(\omega)|^{2}$. One can similarly prove that this transfer function also tends constant in the low-frequency limit and decreases as $\omega^{-2k}$ for some index $k$ in the high-frequency limit.
On the other hand, for a more general case, $\delta a_{1}(t)$ and $\delta a_{-1}(t)$ can be treated as independent, i.e. the cross-correlation $\mathbb{E}[\delta a_{1}(t_{1})\delta a_{-1}(t_{2})]=0$. Then $\delta\phi_{1}$ and $\delta\phi_{-1}$ are also independent, so the variance of dephasing is
$\sigma_{\phi_{\rm diff}}^{2}=\sigma_{\phi_{1}}^{2}+\sigma_{\phi_{-1}}^{2}=\frac{m^{2}}{\hbar^{2}}\int_{0}^{t_{f}}S_{a_{1}a_{1}}(\omega)|x_{1}(\omega)|^{2}+S_{a_{-1}a_{-1}}(\omega)|x_{-1}(\omega)|^{2}\mathrm{d}\omega.$ (35)
As for the Humpty-Dumpty problem, since |$\ket{\beta_{s_{z}}}$ are coherent states and $\beta_{s_{z}}=\beta(t_{f})+\delta\alpha_{s_{z}}$, then the quantity $\braket{\beta_{1}}{\beta_{-1}}$ can be written as
$\braket{\beta_{1}}{\beta_{-1}}=C\mathrm{e}^{i\mathrm{Im}(\beta_{-1}^{*}(t_{f})\beta_{1}(t_{f}))},$ (36)
where $C=\mathrm{e}^{-|\delta\alpha_{1}-\delta\alpha_{-1}|^{2}/2}$ is often refered as the contrast, the Uhlmann fedility [39, 40] or the Loschmidt echo [41, 42, 43] in literature. Therefore, the spatial degrees of freedom can lead to decoherence in the spin space, and this mechanism is known as the Humpty-Dumpty problem [15, 16, 17]. Note that the complex coordinates $\beta_{i}(t_{f})$ in the phase space can be re-written into $x_{i}(t_{f})$ and $p_{i}(t_{f})$ to obtain another form for the contrast loss [15]
$-\log C=\frac{(\Delta x)^{2}}{m\omega_{0}/\hbar}+\frac{(\Delta p)^{2}}{m\hbar\omega_{0}}.$ (37)
Since both $\delta\alpha_{s_{z}}$ follow Gaussian distribution, then $\delta\alpha_{1}-\delta\alpha_{-1}$ also follow Gaussian distribution with a variance $\sigma^{2}_{\rm tot}=\sigma_{1}^{2}+\sigma_{-1}^{2}$ if they are assumed independent. Then the ensemble average of the contrast is
$-\mathbb{E}[\log C]=\frac{1}{2}\mathbb{E}\left[|\delta\alpha_{1}-\delta\alpha_{-1}|^{2}\right]=\frac{\sigma_{1}^{2}+\sigma_{-1}^{2}}{2}.$ (38)
Since $\sigma_{1}$ and $\sigma_{-1}$ are the same, then the contrast loss $-\mathbb{E}[\log C]$ is exactly the variance of a single arm given by (22) which is a linear response to the acceleration noise $\delta a(t)$ with a transfer function $\text{sinc}^{2}(\omega-\omega_{0})t/2$.
To sum up, non-uniform acceleration noises can cause both dephasing and contrast loss on the interferometer, and both effects are linear responses to $\delta a(t)$. Both effects can contribute a decay factor on off-diagonal terms of the spin space, so the observable result for the Ramsey interferometry becomes
$\mathrm{Tr}\left[\mathbb{E}[\hat{\rho}_{\rm spin}(t_{f})]\hat{W}\right]=\frac{1}{2}\left[1+\mathrm{e}^{-\sigma_{\phi_{\rm diff}}^{2}}\mathrm{e}^{-\sigma_{\alpha}^{2}}\cos(\phi_{\rm diff}+\mathrm{Im}(\beta_{-1}^{*}(t_{f})\beta_{1}(t_{f})))\right],$ (39)
A final remark is that by treating some operators as c-numbers, some higher-order effects were not studied. For example, the noise described by $m\delta\omega^{2}(t)\hat{x}^{2}$ like GGN or rotation noise [21] can disorder the shape of the Wigner function by the squeezed operator $\hat{S}(\zeta)=\mathrm{e}^{(\zeta^{*}\hat{a}^{2}-\zeta\hat{a}^{\dagger 2})}$ with $\zeta(t)$ related to $\delta\omega(t)$, then the overlap of the two arms at the final time can be lost even if the center of mass can meet with each other. As for the fluctuation of the magnetic field described by $\mu\delta\eta(t)\hat{S}_{z}\hat{x}$, this noise will affect both the spin space and the phase space, but the decoherence in the spin space cannot be studied by treating $\hat{S}_{z}$ as c-number. Therefore, it is necessary to construct the Lewis-Riesenfeld invariant for these noises to systematically study these higher-order effects in future research.
5 Discussion and Conclusion
5.1 Role of randomness
It is remarkable that the density matrix before and after the ensemble average can be quite different. In particular, the phase fluctuation $\delta\phi_{\rm diff}$ contributes a pure phase factor $\mathrm{e}^{i\delta\phi_{\rm diff}}$ without ensemble average but contributes a decay factor $\mathrm{e}^{-\sigma_{\phi_{\rm diff}}^{2}}$ after average. Similarly, the contrast loss with and without the ensemble average can be also different.
To understand the role of randomness better, one can compute the purity $\text{Tr}[\hat{\rho}^{2}]$ or the von Neumann entropy $S=-\text{Tr}\left[\hat{\rho}\ln\hat{\rho}\right]$. The purities with and without the ensemble average are  [^4] In order to compute the purity and the von Neumann entropy, it is a trick to first compute the eigenvalues of $\hat{\rho}_{\rm spin}$ to obtain the eigenvalues of $\hat{\rho}_{\rm spin}^{2}$ and $\hat{\rho}_{\rm spin}\ln\hat{\rho}_{\rm spin}$. It is not different to check when there is a decay factor $\mathrm{e}^{-\Gamma_{\rm tot}}$ on the off-diagonal terms, the eigenvalues of $\hat{\rho}$ are $\lambda_{1,2}=(1\pm\mathrm{e}^{-\Gamma_{\rm tot}})/2$ and $\lambda_{3}=0$. Then the purity and the entropy are given by $\lambda_{1}^{2}+\lambda_{2}^{2}$ and $\lambda_{1}\ln\lambda_{1}+\lambda_{2}\ln\lambda_{2}$. When the total decoherence factor $\Gamma_{\rm tot}$ is very small, then the entropy can be approximated as $\begin{split}S&=-\left[\frac{1+\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\ln\frac{1+\mathrm{e}^{-\Gamma_{\rm tot}}}{2}+\frac{1-\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\ln\frac{1-\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\right]\approx\frac{1+\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\cdot\frac{1-\mathrm{e}^{-\Gamma_{\rm tot}}}{2}+\frac{1-\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\cdot\frac{1+\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\\ &=\frac{1}{2}(1-\mathrm{e}^{-2\Gamma_{\rm tot}})\approx\Gamma_{\rm tot},\end{split}$ where the first approximated equation used an approximation $\ln((1\pm\mathrm{e}^{-\Gamma_{\rm tot}})/2)\approx(-1\pm\mathrm{e}^{\Gamma_{\rm tot}})/2$.
$\text{Tr}[\hat{\rho}_{\rm spin}^{2}(t_{f})]$ $=\frac{1}{2}(1+C^{2})\approx 1-(-\log C),$ (40)
$\text{Tr}\left[\mathbb{E}[\hat{\rho}_{\rm spin}(t_{f})]^{2}\right]$ $=\frac{1}{2}(1+\mathbb{E}[C]^{2}\mathrm{e}^{-2\sigma_{\phi_{\rm diff}}^{2}})\approx 1-\sigma_{\phi_{\rm diff}}^{2}-\sigma_{\alpha}^{2},$
and the von Neumann entropies with and without the ensemble average are
$S[\hat{\rho}_{\rm spin}]$ $\approx-\log C,$ (41)
$S[\mathbb{E}[\hat{\rho}_{\rm spin}]]$ $\approx\sigma_{\phi_{\rm diff}}^{2}+\sigma_{\alpha}^{2}.$
As is shown from the purity and the entropy, there is only the contrast loss without the ensemble average. When considering the ensemble average, the dephasing effect can cause a purity loss and an entropy increase, and the contrast loss has a different influence on purity and entropy, which can be regarded as a consequence of the position localization decoherence.
Therefore, from the point of view of information, the randomness of the noise plays an essential role in taking out the information of the quantum system, except for the contrast loss mechanism. The information of the quantum system gets lost even if the environment is classical as long as the environment is stochastic.
In summary, the information of the interferometer is lost to the classical noise through three mechanisms. The first mechanism is the dephasing, and the phase information of the two arms is lost due to the randomness of the noise. The second mechanism is the Humpty-Dumpty problem, and the information of the spin space is encoded to the spatial degrees of freedom and gets lost due to the non-uniform noises. The third mechanism is the position localization decoherence, both randomness and non-uniformness of the noise are necessary for this mechanism.
5.2 Summary
In this paper, I did a general theoretical discussion on the decoherence of a Stern-Gerlach interferometer induced by classical stochastic acceleration noises based on the analytic solution derived by the Lewis-Riesenfeld invariant and the quantum Arnold transformation. As is shown in the time-evolution operator, the direct consequences of acceleration noises are the fluctuations of the path integral phase and the classical trajectory in the phase space. Therefore, the first notable conclusion is that the shape of the Wigner function of an arbitrary state doesn’t change but it translates along the classical trajectory and rotates with a constant angular speed of the harmonic oscillator.
For a single arm of the interferometer, the acceleration noise can disturb the path integral phase through two channels in principle. However, as is proved by the equation of motion, the fluctuation of the trajectory and the velocity only contributes a total derivative to the phase, so the phase fluctuation is simply a linear response to the acceleration noise. The fluctuation of the trajectory in the phase space is also a linear response to the noise, and it is a generalized diffusion process of which the long-time limit is approximately a standard Brownian motion. Furthermore, due to the randomness of the noise, the ensemble average of the density matrix can induce a position localization decoherence on the arm of the interferometer.
The witness operator is usually constructed in the spin space, so the degrees of freedom of the phase space have to be traced out. For acceleration noises such as vibrations and inertial forces that are uniform to both arms of the interferometer, it is notable that the classical positions and velocities of both arms at the final time can meet with each other although they are not at the ideal position and velocity, so the witness is only affected by the dephasing effect.
On the other hand, for more generic noises approximately regarded as non-uniform acceleration noises, the final positions in the phase space of the two arms are different, which leads to a contrast loss of the quantum states and finally results in an additional decoherence in the spin space, known as the Humpty-Dumpty problem. In addition, the statistical average of the overlap of the two arms can change the value of the contrast loss, which can be regarded as a consequence of the position localization decoherence.
Finally, it is remarkable that the randomness of the noise plays an essential role in the decoherence. Except for the Humpty-Dumpty problem, the other two mechanisms of the decoherence of the spin space, the dephasing and the position localization decoherence, arise from the ensemble average, which has been illustrated by the purity loss and the von Neumann entropy increase. From the point of view of information theory, the information of the quantum system gets lost due to the randomness of noises through these two mechanisms. Therefore, in an experiment, noises should be monitored as much as possible to eliminate the randomness of the noise, so that the decoherence can be suppressed as much as possible.
Acknowledgements
I would like to thank the China Scholarship Council (CSC) for financial support. I also thank Guan-Zhuo Yang for the beneficial discussions.
Appendix A Analytical Solution for the Schrodinger Equation with Acceleration Noises
In this appendix, I will show the math details to solve the Schrodinger equation with the time-dependent Hamiltonian (2) by the Lewis-Riesenfeld invariant and quantum Arnold transformation method [29, 30, 44, 32, 33].
The basic idea [31] is to find a so-called Lewis-Riesenfeld operator $\hat{I}(t)$ such that all the eigenvalues $\{\lambda_{n}\}$ of $\hat{I}(t)$ are time-independent and all the instantaneous eigenstates {|$\{\ket{n(t)}\}$ of $\hat{I}(t)$ are solutions of Schrodinger equation. Then the time-evolution operator is given by $\hat{U}(t)=\sum_{n}\mathrm{e}^{i\alpha_{n}(t)}\ket{n(t)}\bra{n(0)}$ |, where $\alpha_{n}(t)=\int_{0}^{t}\bra{n(t^{\prime})}\left(i\mathrm{d}/\mathrm{d}t^{\prime}-\hat{H}(t^{\prime})/\hbar\right)\ket{n(t^{\prime})}\mathrm{d}t^{\prime}$ is known as the Lewis-Riesenfeld phase [29, 31]. Under the point view of differential geometry, the instantaneous eigenstates {|$\{\ket{n(t)}\}$ play the role of a moving frame of the Hilbert space as a manifold, and the Lewis-Riesenfeld phase is a kind of geometric phase arising from the connection 1-form of the manifold.
Such time-dependent operator $\hat{I}(t)$ can be generally constructed by a reference frame transformation on the Hamiltonian from the lab reference to the co-moving frame [32], known as the quantum Arnold transformation [33]. In particular, for a driven oscillator with a Hamiltonian (2), consider a translation operator in the phase space of the oscillator $\hat{\mathcal{U}}(t)=\exp\left[-\frac{i}{\hbar}(\hat{x}p(t)-\hat{p}x(t))\right]=\hat{D}(-\alpha(t))$ with $\alpha(t)=\sqrt{m\omega_{0}/2\hbar}(x(t)+ip(t)/m\omega_{0})$. Under the unitary transformation, the wavefunction changes as |$\ket{\psi}\to\ket{\psi^{\prime}}=\hat{\mathcal{U}}(t)\ket{\psi}$, while the Hamiltonian changes as
$\hat{H}(t)\to\hat{\mathcal{U}}\hat{H}\hat{\mathcal{U}}^{\dagger}-i\hbar\hat{\mathcal{U}}\frac{\mathrm{d}}{\mathrm{d}t}\hat{\mathcal{U}}^{\dagger}.$ (42)
The term $-i\hbar\hat{\mathcal{U}}(\mathrm{d}\hat{\mathcal{U}}^{\dagger}/\mathrm{d}t)$ mathematically arises from the connection 1-form to keep the Schrodinger equation covariant under the transformation $\hat{\mathcal{U}}(t)$, and this term is often referred as the conter-diabatic term $\hat{H}_{CD}$ [45]. The calculation details of the transformation of $\hat{H}$ will be summarized in the subsection A.1. The final result is that if the trajectory $x(t)$ and momentum $p(t)$ is chosen satisfying the canonical equation defined by the Hamiltonian (2)
[…]
which is the real-valued form for (4), then the Hamiltonian in the co-moving reference is
$\hat{\mathcal{U}}\hat{H}\hat{\mathcal{U}}^{\dagger}-i\hbar\hat{\mathcal{U}}\frac{\mathrm{d}}{\mathrm{d}t}\hat{\mathcal{U}}^{\dagger}=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}+\delta(t),$ (44)
where the phase $\delta(t)$ is given by
$\delta(t)=-\frac{1}{2}ma(t)x_{c}(t).$ (45)
Therefore, the Hamiltonian in the comoving reference frame is time-independent except for a pure phase term $\delta(t)$. So the time-evolution of the wave function under the comoving reference is simply
|$\ket{\psi^{\prime}(t)}=\mathrm{e}^{-\frac{i}{\hbar}\int_{0}^{t}\delta(t^{\prime})\mathrm{d}t^{\prime}}\mathrm{e}^{-\frac{i}{\hbar}\left(\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}\right)t}\ket{\psi^{\prime}(0)}.$ (46)
Then the wavefunction in the original reference frame is
|$\ket{\psi(t)}=\mathrm{e}^{-\frac{i}{\hbar}\int_{0}^{t}\delta(t^{\prime})\mathrm{d}t^{\prime}}\hat{\mathcal{U}}^{\dagger}(t)\mathrm{e}^{-\frac{i}{\hbar}\left(\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}\right)t}\hat{\mathcal{U}}(0)\ket{\psi(0)}.$ (47)
Furthermore, the equation of motion (43) allows to rewrite the phase $\delta(t)$ in another form
$\delta(t)=-\left(\frac{1}{2}m\dot{x}_{c}^{2}(t)-\frac{1}{2}m\omega_{0}^{2}x_{c}^{2}(t)+ma(t)x_{c}(t)-\frac{1}{2}m\frac{\mathrm{d}}{\mathrm{d}t}(\dot{x}_{c}x_{c})\right).$ (48)
This means that the $\delta(t)$ phase term is exactly the minus Lagrangian $L[x_{c}(t),\dot{x}_{c}(t)]$ of the system (2) along the classical trajectory upto a total derivative term. Thus, the global phase in (47) is exactly the path integral along the classical trajectory $\phi=\int_{0}^{t}L[x_{c}(t^{\prime}),\dot{x}_{c}(t^{\prime})]/\hbar\mathrm{d}t^{\prime}$, so (47) presents the time evolution operator (3). Remarkably, since the $\delta(t)$ phase arises from the gauge term $-i\hbar\hat{\mathcal{U}}(\mathrm{d}\hat{\mathcal{U}}^{\dagger}/\mathrm{d}t)$, the Feynman’s path integral phase in this system has a physical interpretation of geometric phase.
A final remark is that a total derivative term in $\delta(t)$ is usually tolerant in an interferometry experiment because the initial and final position and momentum of different trajectories are the same and they will cancel each other. It is notable that such a total derivative term is related to the choice of the unitary transform operator $\hat{\mathcal{U}}(t)$. In particular, the operator $\hat{\mathcal{U}}(t)$ allows an arbitrary phase and can be chosen as, for example,
$\hat{\mathcal{U}}^{\prime}(t)=\mathrm{e}^{\frac{i}{\hbar}\hat{p}x_{c}(t)}\mathrm{e}^{-\frac{i}{\hbar}\hat{x}p_{c}(t)}=\hat{\mathcal{U}}(t)\mathrm{e}^{\frac{i}{2\hbar}p_{c}(t)x_{c}(t)}.$ (49)
One can use $\hat{\mathcal{U}}^{\prime}(t)$ to do the same calculation of $\hat{\mathcal{U}}^{\prime}\hat{H}\hat{\mathcal{U}}^{\prime\dagger}-i\hbar\hat{\mathcal{U}}^{\prime}(\mathrm{d}\hat{\mathcal{U}}^{\prime\dagger}/\mathrm{d}t)$, and find the additional phase term $\mathrm{e}^{(i/2\hbar)p_{c}(t)x_{c}(t)}$ in $\hat{\mathcal{U}}^{\prime}$ will contribute a c-number term $1/2(\dot{p}_{c}(t)x_{c}(t)+\dot{x}_{c}(t)p_{c}(t))$ in $-i\hbar\hat{\mathcal{U}}(\mathrm{d}\hat{\mathcal{U}}^{\prime\dagger}/\mathrm{d}t)$. Then the last term in (56) becomes $x_{c}(t)\dot{p}_{c}(t)$ rather than $1/2(x_{c}(t)\dot{p}_{c}(t)-p_{c}(t)\dot{x}_{c}(t))$, and the corresponding pure phase term $\delta(t)$ becomes [46]
$\delta^{\prime}(t)=\frac{1}{2}m\dot{x}_{c}^{2}(t)-\frac{1}{2}m\omega_{0}^{2}x_{c}^{2}(t).$ (50)
Based on the equation of motion (43), one can easily check the equivalence between $\delta(t)$ and $\delta^{\prime}(t)$ upto a total derivative term $\frac{1}{2}m\frac{\mathrm{d}}{\mathrm{d}t}(x_{c}\dot{x}_{c})$, i.e.
$-\frac{1}{2}ma(t)x_{c}(t)=\frac{1}{2}m\dot{x}_{c}^{2}(t)-\frac{1}{2}m\omega_{0}^{2}x_{c}^{2}(t)-\frac{1}{2}m\frac{\mathrm{d}}{\mathrm{d}t}(x_{c}\dot{x}_{c}).$ (51)
A.1 Calculation details of the transformation of Hamiltonian
The calculation details to compute the transformation of the Hamiltonian under the quantum Arnold transformation (42) is summarized as follows.
To start with, recall the Campbell formula $\mathrm{e}^{\hat{A}}\hat{B}\mathrm{e}^{-\hat{A}}=\hat{B}+[\hat{A},\hat{B}]$ if $[\hat{A},[\hat{A},\hat{B}]]=0$. Then for the reference translation operator $\hat{\mathcal{U}}(t)=\exp\left[-\frac{i}{\hbar}(\hat{x}p(t)-\hat{p}x(t))\right]$, one can compute that
$\hat{\mathcal{U}}\hat{x}\hat{\mathcal{U}}^{\dagger}$ $=\hat{x}+x_{c}(t),$ (52)
$\hat{\mathcal{U}}\hat{p}\hat{\mathcal{U}}^{\dagger}$ $=\hat{p}+p_{c}(t).$
Note that a unitary transformation of any arbitrary analytic function of any arbitrary operator $f(\hat{X})$ satisfies $\hat{\mathcal{U}}f(\hat{X})\hat{\mathcal{U}}^{\dagger}=f(\hat{\mathcal{U}}\hat{X}\hat{\mathcal{U}}^{\dagger})$, then the transformation $\hat{\mathcal{U}}\hat{H}(t)\hat{\mathcal{U}}^{\dagger}$ can be computed as
$\begin{split}\hat{\mathcal{U}}\hat{H}(t)\hat{\mathcal{U}}^{\dagger}&=\frac{\left(\hat{p}+p_{c}(t)\right)^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\left(\hat{x}+x_{c}(t)\right)^{2}-ma(t)\left(\hat{x}+x_{c}(t)\right)\\ &=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}+\frac{p_{c}(t)}{m}\hat{p}+(m\omega_{0}^{2}x_{c}(t)-ma(t))\hat{x}+\frac{p_{c}^{2}(t)}{2m}+\frac{1}{2}m\omega_{0}^{2}x_{c}^{2}(t)-ma(t)x_{c}(t).\end{split}$ (53)
For the gauge potential term, one needs to first compute $\hat{\mathcal{U}}^{\dagger}(t+\mathrm{d}t)$, i.e.
$\begin{split}\hat{\mathcal{U}}^{\dagger}(t+\mathrm{d}t)&=\exp\left[\frac{i}{\hbar}\left(\hat{x}p_{c}(t+\mathrm{d}t)-\hat{p}x_{c}(t+\mathrm{d}t)\right)\right]\\ &=\exp\left[\frac{i}{\hbar}(\hat{x}p_{c}(t)-\hat{p}x_{c}(t))\right]\exp\left[\frac{i}{\hbar}(\hat{x}\dot{p}_{c}(t)-\hat{p}\dot{x}_{c}(t))\mathrm{d}t\right]\exp\left[-\frac{i}{2\hbar}(x_{c}(t)\dot{p}_{c}(t)-p_{c}(t)\dot{x}_{c}(t))\mathrm{d}t\right]\\ &=\hat{\mathcal{U}}^{\dagger}(t)\left(1+\frac{i}{\hbar}(\hat{x}\dot{p}_{c}(t)-\hat{p}\dot{x}_{c}(t))\mathrm{d}t+\mathcal{O}((\mathrm{d}t)^{2})\right)\left(1+\frac{i}{2\hbar}(x_{c}(t)\dot{p}_{c}(t)-p_{c}(t)\dot{x}_{c}(t))\mathrm{d}t+\mathcal{O}((\mathrm{d}t)^{2})\right)\\ &=\hat{\mathcal{U}}^{\dagger}(t)\left(1+\frac{i}{\hbar}(\hat{x}\dot{p}_{c}(t)-\hat{p}\dot{x}_{c}(t))\mathrm{d}t+\frac{i}{2\hbar}(x_{c}(t)\dot{p}_{c}(t)-p_{c}(t)\dot{x}_{c}(t))\mathrm{d}t+\mathcal{O}((\mathrm{d}t)^{2})\right).\end{split}$ (54)
Then the gauge potential $-i\hbar\hat{\mathcal{U}}\mathrm{d}\hat{\mathcal{U}}^{\dagger}/\mathrm{d}t$ is
$-i\hbar\hat{\mathcal{U}}(t)\frac{\mathrm{d}}{\mathrm{d}t}\hat{\mathcal{U}}^{\dagger}(t)=\hat{x}\dot{p}_{c}(t)-\hat{p}\dot{x}_{c}(t)+\frac{1}{2}\left(x_{c}(t)\dot{p}_{c}(t)-p_{c}(t)\dot{x}_{c}(t)\right).$ (55)
Therefore, the Hamiltonian in the comoving reference frame, i.e. under the transformation $\hat{\mathcal{U}}(t)$ is
$\begin{split}\hat{\mathcal{U}}(t)\hat{H}(t)\hat{\mathcal{U}}^{\dagger}(t)-i\hbar\hat{\mathcal{U}}(t)\frac{\mathrm{d}}{\mathrm{d}t}\hat{\mathcal{U}}^{\dagger}(t)&=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}+\left(\frac{p_{c}(t)}{m}-\dot{x}_{c}(t)\right)\hat{p}+\left(m\omega_{0}^{2}x_{c}(t)+\dot{p}_{c}(t)-ma(t)\right)\hat{x}\\ &+\frac{p_{c}^{2}(t)}{2m}+\frac{1}{2}m\omega_{0}^{2}x_{c}^{2}(t)-ma(t)x_{c}(t)+\frac{1}{2}\left(x_{c}(t)\dot{p}_{c}(t)-p_{c}(t)\dot{x}_{c}(t)\right).\end{split}$ (56)
If one chooses the trajectory $x_{c}(t)$ and momentum $p_{c}(t)$ as the classical trajectory and momentum defined by the Hamiltonian (2) satisfying the canonical equation of motion (43), then the Hamiltonian in the comoving reference frame can be simplified as
$\hat{\mathcal{U}}\hat{H}(t)\hat{\mathcal{U}}^{\dagger}-i\hbar\hat{\mathcal{U}}\frac{\mathrm{d}}{\mathrm{d}t}\hat{\mathcal{U}}^{\dagger}=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}+\delta(t),$ (57)
where the pure phase term $\delta(t)$ is given by (45). As is pointed out in (48), $\delta(t)$ is exactly the minus of the Lagrangian of the system, so the related phase is exactly Feynman’s path integral phase.
A.2 Time-evolution of Wigner function
From the point of view of Weyl quantization, the Wigner function is the Weyl-Wigner transform of the density matrix, given by
$W(\alpha)=\frac{1}{\pi^{2}}\int\text{Tr}[\hat{\rho}\hat{D}(\alpha^{\prime})]\mathrm{e}^{\alpha^{\prime*}\alpha-\alpha^{\prime}\alpha^{*}}\mathrm{d}^{2}\alpha^{\prime}.$ (58)
Since the time-evolution of the density matrix is $\hat{\rho}(t)=\hat{U}(t)\hat{\rho}(0)\hat{U}^{\dagger}(t)$, where the time evolution operator $\hat{U}(t)$ is given by (3), then the trace inside the integral of the Wigner function can be computed as
$\begin{split}\text{Tr}[\hat{\rho}(t)\hat{D}(\alpha^{\prime})]&=\text{Tr}[\hat{\rho}(0)\hat{D}^{\dagger}(-\alpha_{c}(0))\mathrm{e}^{\frac{i}{\hbar}\hat{H}_{0}t}\hat{D}^{\dagger}(\alpha_{c}(t))\hat{D}(\alpha^{\prime})\hat{D}(\alpha_{c}(t))\mathrm{e}^{-\frac{i}{\hbar}\hat{H}_{0}t}\hat{D}(-\alpha_{c}(0))]\\ &=\text{Tr}[\hat{\rho}(0)\hat{D}^{\dagger}(-\alpha_{c}(0))\mathrm{e}^{\frac{i}{\hbar}\hat{H}_{0}t}\hat{D}(\alpha^{\prime})\mathrm{e}^{-\frac{i}{\hbar}\hat{H}_{0}t}\hat{D}(-\alpha_{c}(0))]\mathrm{e}^{\alpha^{\prime}\alpha_{c}^{*}(t)-\alpha^{\prime*}\alpha_{c}(t)}\\ &=\text{Tr}[\hat{\rho}(0)\hat{D}^{\dagger}(-\alpha_{c}(0))\hat{D}(\alpha^{\prime}\mathrm{e}^{i\omega_{0}t})\hat{D}(-\alpha_{c}(0))]\mathrm{e}^{\alpha^{\prime}\alpha_{c}^{*}(t)-\alpha^{\prime*}\alpha_{c}(t)}\\ &=\text{Tr}[\hat{\rho}(0)\hat{D}(\alpha^{\prime}\mathrm{e}^{i\omega_{0}t})]\mathrm{e}^{\alpha^{\prime}\alpha_{c}^{*}(t)-\alpha^{\prime*}\alpha_{c}(t)}\mathrm{e}^{\alpha^{\prime}\mathrm{e}^{i\omega_{0}t}(-\alpha_{c}^{*}(0))-\alpha^{\prime*}\mathrm{e}^{-i\omega_{0}t}(-\alpha_{c}(0))}\\ &=\text{Tr}[\hat{\rho}(0)\hat{D}(\alpha^{\prime}\mathrm{e}^{i\omega_{0}t})]\mathrm{e}^{(\alpha^{\prime}\mathrm{e}^{i\omega_{0}t})^{*}(\alpha_{c}(0)-\alpha_{c}(t)\mathrm{e}^{i\omega_{0}t})-(\alpha^{\prime}\mathrm{e}^{i\omega_{0}t})(\alpha_{c}(0)-\alpha_{c}(t)\mathrm{e}^{i\omega_{0}t})^{*}}\\ &=\text{Tr}[\hat{\rho}(0)\hat{D}(\alpha^{\prime\prime})]\mathrm{e}^{\alpha^{\prime\prime*}(\alpha_{c}(0)-\alpha_{c}(t)\mathrm{e}^{i\omega_{0}t})-c.c.},\end{split}$ (59)
where $\alpha^{\prime\prime}=\alpha^{\prime}\mathrm{e}^{i\omega_{0}t}$ and $c.c.$ stands for the complex conjugate. Therefore, by a variable substitution, one can compute the time-evolution of the Wigner function as
$\begin{split}W(\alpha,t)&=\frac{1}{\pi^{2}}\int\text{Tr}[\hat{\rho}(t)\hat{D}(\alpha^{\prime})]\mathrm{e}^{\alpha^{\prime*}\alpha-\alpha^{\prime}\alpha^{*}}\mathrm{d}^{2}\alpha^{\prime}\\ &=\frac{1}{\pi^{2}}\int\text{Tr}[\hat{\rho}(0)\hat{D}(\alpha^{\prime\prime})]\mathrm{e}^{\alpha^{\prime\prime*}(\alpha_{c}(0)-\alpha_{c}(t)\mathrm{e}^{i\omega_{0}t}+\alpha\mathrm{e}^{i\omega_{0}t})-c.c.}\mathrm{d}\alpha^{\prime\prime}\\ &=W(\alpha_{c}(0)+\mathrm{e}^{i\omega_{0}t}(\alpha-\alpha_{c}(t)),t=0).\end{split}$ (60)
Appendix B Witness for Ramsey Interferometry
[…]
$\hat{H}_{p}=\hbar\Omega_{p}\begin{pmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{pmatrix}=\hbar\Omega_{p}\hat{T}^{\dagger}\begin{pmatrix}\sqrt{2}&0&0\\ 0&0&0\\ 0&0&-\sqrt{2}\end{pmatrix}\hat{T},$ (61)
where the transform matrix $\hat{T}$ is
$\hat{T}=\begin{pmatrix}\frac{1}{2}&\frac{1}{\sqrt{2}}&\frac{1}{2}\\ \frac{1}{\sqrt{2}}&0&-\frac{1}{\sqrt{2}}\\ \frac{1}{2}&-\frac{1}{\sqrt{2}}&\frac{1}{2}\end{pmatrix}.$ (62)
Then the time-evolution operator $\hat{U}_{p}(t)=\mathrm{e}^{-i\hat{H}_{p}t/\hbar}$ for the pulse acting on the spin-1 system is given by
$\hat{U}_{p}(t)=\hat{T}^{\dagger}\begin{pmatrix}\mathrm{e}^{-i\sqrt{2}\Omega_{p}t}&0&0\\ 0&1&0\\ 0&0&\mathrm{e}^{i\sqrt{2}\Omega_{p}t}\end{pmatrix}\hat{T}=\begin{pmatrix}\cos^{2}\frac{\sqrt{2}\Omega_{p}t}{2}&-\frac{i}{\sqrt{2}}\sin\sqrt{2}\Omega_{p}t&-\sin^{2}\frac{\sqrt{2}\Omega_{p}t}{2}\\ -\frac{i}{\sqrt{2}}\sin\sqrt{2}\Omega_{p}t&\cos\sqrt{2}\Omega_{p}t&-\frac{i}{\sqrt{2}}\sin\sqrt{2}\Omega_{p}t\\ -\sin^{2}\frac{\sqrt{2}\Omega_{p}t}{2}&-\frac{i}{\sqrt{2}}\sin\sqrt{2}\Omega_{p}t&\cos^{2}\frac{\sqrt{2}\Omega_{p}t}{2}\end{pmatrix}.$ (63)
If the time of the microwave pulse is set as $t_{p}=\pi/(2\sqrt{2}\Omega_{p})$, then the corresponding time-evolution operator is
$\hat{U}_{\pi/2}\equiv\hat{U}_{p}\left(t_{p}=\frac{\pi}{2\sqrt{2}\Omega_{p}}\right)=\begin{pmatrix}\frac{1}{2}&-\frac{i}{\sqrt{2}}&-\frac{1}{2}\\ -\frac{i}{\sqrt{2}}&0&-\frac{i}{\sqrt{2}}\\ -\frac{1}{2}&-\frac{i}{\sqrt{2}}&\frac{1}{2}\end{pmatrix}.$ (64)
As two examples, $\pi/2$-pulse makes the initial state |$\ket{0}$ to a state $\hat{U}_{\pi/2}\ket{0}=\mathrm{e}^{-i\frac{\pi}{2}}(\ket{1}+\ket{-1})/\sqrt{2}$, and for the state |$\ket{\psi(t_{f})}=(\ket{1}+\mathrm{e}^{i\phi_{\rm diff}}\ket{-1})/\sqrt{2}$, the $\pi/2$-pulse makes it to a state
$\hat{U}_{\pi/2}\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ 0\\ \mathrm{e}^{i\phi_{\rm diff}}\end{pmatrix}=\frac{1}{\sqrt{2}}\mathrm{e}^{-i\frac{\pi}{2}}\mathrm{e}^{i\frac{\phi_{\rm diff}}{2}}\begin{pmatrix}\sin\frac{\phi_{\rm diff}}{2}\\ \sqrt{2}\cos\frac{\phi_{\rm diff}}{2}\\ -\sin\frac{\phi_{\rm diff}}{2}\end{pmatrix}.$ (65)
After the $\pi/2$-pulse, the witness to measure the population of |$\ket{S_{z}=0}$ is simply $\hat{W}^{\prime}=\ket{S_{z}=0}\bra{S_{z}=0}$ |. Since the $\pi/2$-pulse makes a transformation on $\hat{\rho}_{\rm spin}$ to $\hat{U}_{\pi/2}\hat{\rho}_{\rm spin}\hat{U}_{\pi/2}^{\dagger}$, the observable result for the witness $\hat{W}^{\prime}$ is $\mathrm{Tr}[\hat{W}^{\prime}\hat{U}_{\pi/2}\hat{\rho}_{\rm spin}\hat{U}_{\pi/2}^{\dagger}]$. Therefore, the witness for Ramsey interferometry is given by
$\hat{W}=\hat{U}_{\pi/2}^{\dagger}\hat{W}^{\prime}\hat{U}_{\pi/2}=\begin{pmatrix}\frac{1}{2}&0&\frac{1}{2}\\ 0&0&0\\ \frac{1}{2}&0&\frac{1}{2}\end{pmatrix}.$ (66)
Appendix C The Master Equation of the Decoherence of a Single Path
Let’s work under the interaction picture to derive the master equation for the decoherence of a single path. To start with, the Hamiltonian (2) of a single arm can be divided into $\hat{H}_{0}$ and $\hat{H}_{\rm int}(t)=-F(t)\hat{x}$, where $\hat{H}_{0}$ is exactly the Hamiltonian of a simple oscillator and all the force terms are included in the latter term. So the driven force and the signal will perform as the mean value of $F(t)$, while the acceleration noise affects the auto-correlation of $F(t)$. Then the dynamical equation for $\hat{\rho}^{I}$ is
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{\rho}^{I}=-\frac{i}{\hbar}[\hat{H}^{I}_{\rm int}(t),\hat{\rho}^{I}(t)].$ (67)
This differential equation can be directly written in an integral form as
$\hat{\rho}^{I}(t)=\hat{\rho}^{I}(0)-\frac{i}{\hbar}\int_{0}^{t}[\hat{H}^{I}_{\rm int}(t^{\prime}),\hat{\rho}^{I}(t^{\prime})]\mathrm{d}t^{\prime}.$ (68)
The idea to derive the master equation is to do recursion for $\hat{\rho}^{I}(t)$, that is, to insert (68) to the right side of (67), then one may obtain
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\hat{\rho}^{I}&=-\frac{i}{\hbar}[\hat{H}^{I}_{\rm int}(t),\hat{\rho}^{I}(0)-\frac{i}{\hbar}\int_{0}^{t}[\hat{H}^{I}_{\rm int}(t^{\prime}),\hat{\rho}^{I}(t^{\prime})]\mathrm{d}t^{\prime}]\\ &=-\frac{1}{\hbar^{2}}\int_{0}^{t}[\hat{H}^{I}_{\rm int}(t),[\hat{H}^{I}_{\rm int}(t^{\prime}),\hat{\rho}^{I}(t^{\prime})]]\mathrm{d}t^{\prime},\end{split}$ (69)
where $\hat{\rho}^{I}(0)$ is assumed to commute with $\hat{H}^{I}_{\rm int}(t)$, i.e. $[\hat{H}^{I}_{\rm int}(t),\hat{\rho}^{I}(0)]=0$. Then the decoherence of the density matrix will arise from the ensemble average of the total driven force $F(t)$ as
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[\hat{\rho}^{I}(t)]=-\frac{1}{\hbar^{2}}\int_{0}^{t}\mathbb{E}[f(t)f(t^{\prime})][\hat{x}^{I}(t),[\hat{x}^{I}(t^{\prime}),\hat{\rho}^{I}(t^{\prime})]]\mathrm{d}t^{\prime}.$ (70)
This is a deformation of the Redfield master equation [47] for a classical stochastic noise. Furthermore, the decoherence is usually assumed memoryless, known as the Markovian assumption [26, 47] which means that the $\hat{\rho}^{I}(t^{\prime})$ in (70) can be assumed to only depend on $t$, so the master equation is
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[\hat{\rho}^{I}(t)]=-\frac{1}{\hbar^{2}}[\hat{x}^{I},[\hat{x}^{I},\mathbb{E}[\hat{\rho}^{I}(t)]]]\int_{0}^{t}\mathbb{E}[f(t)f(t^{\prime})]\mathrm{d}t^{\prime}.$ (71)
By switching the interaction picture to the Schrodinger picture, one may obtain the master equation (27) in the main text with $\Lambda=\int_{0}^{t}\mathbb{E}[f(t)f(t^{\prime})]/\hbar^{2}\mathrm{d}t^{\prime}$.
Reference