### v3

# Acceleration Noise Induced Decoherence in Stern-Gerlach Interferometers for Gravity Experiments

Abstract
Stern-Gerlach interferometer (SGI) is a kind of matter-wave interferometer driven by magnetic field and has been proposed for various gravity experiments. Stochastic noises can lead to decoherence problems of SGI via several typical mechanisms including dephasing, loss of contrast (known as the Humpty-Dumpty problem) and position localisation decoherence. This paper aims to theoretically explore these mechanisms by modelling the test mass as a harmonic oscillator under acceleration noises described by Gaussian stochastic processes. As will be demonstrated, the dephasing is a linear response to the noise, with the transfer function as the Fourier transform of the unperturbed classical trajectories. Besides, the noise-induced fidelity loss and spatial decoherence of each single path will cancel for common mode noises. By contrast, path-dependent noises can lead to a loss of Loschmidt echo and spatial decoherence characterised by the PSD of the noise, resulting in a loss of witness constructed in spin space. Remarkably, the randomness of the noise is crucial for dephasing and position-localisation decoherence. These two mechanisms don’t cause purity loss or entropy increase if the noise is determinstic rather than stochastic.
1 Introduction
Stern-Gerlach interferometer (SGI) is a type of matter-wave interferometer driven by an external magnetic field through the spin of the system [1]. In recent years, SGIs based on Nitrogen-Vacancy (NV) centres [2] has been proposed as gravity sensors [3] like gravimeters [4, 5] and gyroscopes [6, 7]. They have also been proposed for exploring the nature of gravity [8]. Compared to the matter-wave interferometers driven by lasers, the test mass of an SGI is usually much heavier. Consequently, the amplitude of the signal of gravitational interactions can be significantly larger and easier to detect.
Robustness is an essential issue for all types of matter-wave interferometers, including the SGI. In real experiments, various sources of classical acceleration noises arise from mechanical vibrations, seismic motions, inertial forces and more, while other noises couple to the test mass in different ways, including gravity gradient (tidal force) noise, rotation noises and magnetic field noise. These noises can disturb the trajectories and velocities of the two arms of the interferometer at the classical level and can cause decoherence of the quantum state. Since noises can be characterised by their power spectrum densities (PSD), a fundamental question arises: how to formulate the decoherence factors caused by noises based on their PSDs? Numerous previous studies have investigated the decoherence of SGIs, classifying it into several typical mechanisms including the dephasing, the position localisation decoherence [9, 10, 11, 12, 13, 14] and the loss of contrast (known as the Humpty-Dumpty problem[15, 16, 17].
The dephasing effect, which arises from the ensemble average of random phase fluctuations induced by stochastic noise, can be classified into the phase fluctuation of both the spin phase [18, 19, 20] and the spatial path integral phase [21, 22, 23, 24]. In spin space, dephasing is typically characterised by the $T_{2}$ time in the Bloch-Redfield equation, which is a linear response to the PSD of the noise [19, 20]. As for spatial dephasing, some researchers have observed that it is also a linear response to the noise, characterised by a transfer function equal to the Fourier transform of the classical unperturbed trajectories [21, 22, 23, 24]. Nonetheless, several subtle gaps remain: why does the dephasing depend solely on the classical trajectories of the centre of mass; and why does the fluctuation of the trajectories provide no contribution to the dephasing?
Position-localisation decoherence, which pertains to the decoherence of the spatial wavefunctions, was originally studied to elucidate the collapse model [9] of quantum states caused by the interactions with the ambient environment. This type of decoherence is typically attributed to scattering processes involving particles such as air molecules or photons [9, 25, 10]. Furthermore, the ensemble average of stochastic noises can also induce a dissipator $\mathcal{D}[\cdot]=-\Lambda[\hat{x},[\hat{x},\cdot]]$, leading to a decay factor $\mathrm{e}^{-\Lambda(x_{1}-x_{2})^{2}t}$ applied to the spatial density matrix [26, 27, 12, 14]. In a scenario involving SGI, the witness operator is usually constructed within the spin space, prompting a question how the position localisation decoherence affects the spin witness.
Both the dephasing effect and the position-localisation decoherence occur in all types of interferometers, but the Humpty-Dumpty problem is a peculiar issue in SGI, especially in full-loop SGI [28, 1]. In this type of interferometers, the witness is usually constructed in the spin space, necessitating to trace out of spatial degrees of freedom. Consequently, the loss of the Loschmit echo [29] (also referred to as the contrast) can lead to a decay factor of the off-diagonal terms in the spin density matrix, known as the Humpty-Dumpty problem. The most fundamental source of the contrast loss arises from the non-closure of the two arms of the interferometer at the final time [16, 30, 28]. Other plausible sources have been theoretically studied in recent researches, such as the rotation effect [31], the coupling with phonons [32] and the deformation of the wave-packet [33]. Notably, classical noises can also cause contrast loss problem, warranting further investigation to formulate the associated contrast loss.
Moreover, it is evident that all these mechanisms are related to the PSD of the noise, but previous papers studied them separately. Therefore, this underscores the need for systematic research that incorporates all these mechanisms via the full dynamics of the test mass, despite mathematically challenging posed by the time dependence and randomness of noises. Consequently, most previous researches involving the full dynamics have focused on the white noise [5, 34, 35], which is identified by a constant PSD and lacks time correlation. Some authors [36] have noticed the limitations of analysing white noise, prompting them to consider the time dependence of the noise and to solve the corresponding Schrödinger equation using perturbation methods. However, the connection between the decoherence factor and the PSD remains unestablished.
In light of the aforementioned issues, this paper aims to formulate the decoherence mechanisms outlined above based on the PSD of the external noise. The system will be modelled as a harmonic oscillator under a driven force with fluctuations, and its time evolution can be analytically solved using an elegant method known as the Lewis-Resenfeld invariant [37, 38, 39, 40] and quantum Arnold transformation [41].
Building upon the analytical solution, I will first conduct a systematic discussion on the dephasing as a linear response, particularly focusing on validating the subtle propositions mentioned above. Besides, I will also investigate the impact of position localisation decoherence on the spin witness. The central question of this paper is to explore the contrast loss arising from non-closure problem due to classical noises. In particular, I will first demonstrate the invariance of the wavefunction shape in the presence of acceleration noise fluctuations. Then I will explore the fluctuation of the classical trajectories, and find out the resulting fidelity loss in a single arm and the overall contrast loss at the final time.
This paper is organised as follows. In section 2, the dynamics of an ideal SGI without noises will be briefly reviewed, with sub-section 2.1 studying the motion of the spatial part and sub-section 2.2 discussing the witness constructed in the spin space. The mathematical details for sub-sections 2.1 and 2.2 are summarized in Appendix A and B in respect. In section 3, the influence of acceleration noises on a single arm will be firstly studied, including the dephasing, the fluctuation of the classical trajectory in the phase space, the fidelity loss and the position localisation decoherence, which will be discussed separately in sub-sections. The master equation of the position localisation decoherence of a single arm will be derived in Appendix C. Sub-section 3.4 will study the decoherence in the spin space and the corresponding loss of witness of the interferometer. Section 4 will make some discussions, with sub-section 4.1 addressing the decoherence caused by path-dependent noises, sub-section 4.2 discussing the role of randomness of the noise, and sub-section 4.3 providing a summary of this paper.
2 Ideal Stern-Gerlach Interferometer without Noises
$t$$x$$B(t)=\eta(t)x$$mg$|$\ket{S}=(\ket{\uparrow}+\ket{\downarrow})/\sqrt{2}$|$\ket{S}=(\ket{\uparrow}+\mathrm{e}^{i\phi_{\rm diff}}\ket{\downarrow})/\sqrt{2}$
Figure 1: The schematic diagram of an ideal Stern-Gerlach interferometer for gravity experiments. Under a non-uniform magnetic field $\mathbf{B}=\eta\mathbf{x}$, the test mass splits to two trajectories with respect to different spin states. The magnetic field gradient $\eta\equiv\partial B/\partial x$ can be set as a time-independent constant [4], a piecewise function of time [36], several pulses [28], or more complicated functions [42, 33]. The differential phase $\phi_{\rm diff}$ at the final time encodes the information of gravity, so properties of gravity can be detected by measuring this phase.
A general schematic diagram of a Stern-Gerlach interferometer is shown as Fig. 1. The test mass is trapped in a potential well which can be assumed as a harmonic oscillator for simplicity. Besides, the spin of the test mass has to be splitted during the experiment, which can be achieved by some intrinsic properties of the test mass or a constant biased external magnetic field (i.e. the Zeeman effect). Then the Stern-Gerlach interferometer can be created by a non-uniform external magnetic field. For example, the 1-dimensional Hamiltonian of the ideal gravimeter without noise based on the nitrogen-vacancy (NV) centre can be written as [36]
$\hat{H}=\frac{\hat{p}^{2}}{2m}+g_{\rm NV}\mu_{B}\frac{\partial B}{\partial x}\hat{S}_{z}\hat{x}-\frac{\chi V}{2\mu_{0}}\left(\frac{\partial B}{\partial x}\right)^{2}\hat{x}^{2}-mg\hat{x}+\hbar D\hat{S}_{z}^{2}.$ (1)
The term $\hbar D\hat{S}_{z}^{2}$ is the zero-field splitting of the NV centre with $D=(2\pi)2.8$ GHz, which physically arises from the electron spin-spin dipole interaction [43]. Under this term, the spin of the test mass splits to eigenstates of $\hat{S}_{z}$, where $z$-direction is assumed as the direction of the NV-axis for simplicity, which can be changed to any axis in a real experiment. The spin-orbit coupling term $g_{\rm NV}\mu_{B}(\partial B/\partial x)\hat{S}_{z}\hat{x}$ arises the spin-magnetic field interaction, where $\mu_{B}$ is the Bohr magneton and $g_{\rm NV}$ is the Lande g-factor of NV centre. The quadratic term $\chi V/(2\mu_{0})(\partial B/\partial x)^{2}\hat{x}^{2}$ is the diamagnetic energy of the diamond, with $\chi=-2.2\times 10^{-5}$ and $V$ are the magnetic susceptibility and the volume of the diamond. The magnetic field gradient $\partial B/\partial x$ can be set as a time-independent constant [4], a piecewise function of time [36], several pulses [28], or more complicated functions [42, 33]. The term $mg\hat{x}$ accounts for the coupling to gravitational acceleration, which can be substituted with other coupling mechanisms for different experimental contexts. For instance, this sensor can also function as an accelerometer or seismometer by coupling to inertial forces [3, 44].
In order to create superpositions by the magnetic field, it is essential to avoid the procession of the spin such that spin states are pinned along the direction defined by the splitting term. For the Hamiltonian (1), the $\hat{\mathbf{S}}\cdot\mathbf{B}$ term is supposed smaller than the zero-splitting term to avoid the Larmor precession. Since the energy of the zero-splitting term is $(2\pi)\hbar D\sim 1.8\times 10^{-24}\,\textrm{J}\sim g\mu_{B}\times 10^{-1}\,\textrm{T}$, then the value of the magnetic field $|\mathbf{B}|$ is required smaller than 0.1 T [31]. In this case, the spin operator $\hat{S}_{z}$ can be treated as a c-number rather than an operator for ease of calculation. Then $s_{z}=\pm 1$ of the interferometer corresponds to two paths, and each path of the interferometer is exactly a trajectory of a harmonic oscillator driven by external forces. Thus, the Hamiltonian for a single arm can be written as
$\hat{H}(t)=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}-ma(t)\hat{x},$ (2)
where $ma(t)$ consists of the driven force $g_{\rm NV}\mu_{B}(\partial B/\partial x)s_{z}\hat{x}$ and the signal force $mg\hat{x}$, where the time-dependence of the Hamiltonian mainly comes from the time-dependent noise. The corresponding Schrodinger equation can be analytically solved by Lewis-Riesenfeld invariant and quantum Arnold transformation [40, 41], and the math details are summarized in A. Notably, other interferometers driven by spin-orbit coupling (SOC) [45] or SQUID [46] have Hamiltonians of similar forms, so the noise analysis in this paper is also valid for these types of interferometers.
2.1 Time evolution of single path
As is derived in A, the time-evolution operator for the Hamiltonian (2) is
$\hat{U}(t)=\mathrm{e}^{\frac{i}{\hbar}\int_{0}^{t}L[x(t^{\prime}),\dot{x}(t^{\prime})]\mathrm{d}t^{\prime}}\hat{D}(\alpha_{c}(t))\mathrm{e}^{-\frac{i}{\hbar}\hat{H}_{0}t}\hat{D}(-\alpha_{c}(0)),$ (3)
where $L[x(t),\dot{x}(t)]=\frac{1}{2}m\dot{x}^{2}-\frac{1}{2}m\omega_{0}^{2}x+ma(t)x$ is the Lagrangian corresponding to the Hamiltonian (2), $\hat{H}_{0}=\hat{p}^{2}/2m+m\omega_{0}^{2}\hat{x}^{2}/2$ is the Hamiltonian of a simple harmonic oscillator without driven forces, and $\hat{D}(\alpha_{c}(t))=\mathrm{e}^{\alpha_{c}(t)\hat{a}^{\dagger}-\alpha_{c}^{*}(t)\hat{a}}$ is the displacement operator defined by the complex coordinate $\alpha_{c}(t)=\sqrt{m\omega_{0}/2\hbar}(x_{c}(t)+ip_{c}(t)/m\omega_{0})$ of the instantaneous equilibrium position in the phase space, which satisfies the canonical equation in complex coordinate
$\dot{\alpha}_{c}=-i\omega_{0}\alpha_{c}+i\sqrt{\frac{m}{2\hbar\omega_{0}}}a(t),$ (4)
of which the solution is
$\alpha_{c}(t)=\mathrm{e}^{-i\omega_{0}t}\left(\alpha_{c}(0)+i\sqrt{\frac{m}{2\hbar\omega_{0}}}\int_{0}^{t}a(t^{\prime})\mathrm{e}^{i\omega_{0}t^{\prime}}\mathrm{d}t^{\prime}\right).$ (5)
The physical interpretation of the time evolution operator $\hat{U}(t)$ in (3) is that the system is a simple oscillator without any driven force in the co-moving reference frame of the equilibrium position, where the displacement operators $\hat{D}(\alpha_{c}(t))$ and $\hat{D}(-\alpha_{c}(0))$ play the role of the reference translation in the classical phase space. Besides, $\hat{U}(t)$ also contains a path integral phase $\mathrm{e}^{\frac{i}{\hbar}\int_{0}^{t}L[x(t^{\prime}),\dot{x}(t^{\prime})]\mathrm{d}t^{\prime}}$ along the classical trajectory of the equilibrium position $x_{c}(t)$ of the oscillator.
Consider an arbitrary initial state described by a density matrix $\hat{\rho}_{0}$ evolving as $\hat{\rho}(t)=\hat{U}(t)\hat{\rho}_{0}\hat{U}^{\dagger}(t)$. From the point of view of Weyl quantization, the Wigner function is the Weyl-Wigner transform of the density matrix [47], i.e. $W(\alpha)=\int\text{Tr}[\hat{\rho}\hat{D}(\alpha^{\prime})]\mathrm{e}^{\alpha^{\prime*}\alpha-\alpha^{\prime}\alpha^{*}}/\pi^{2}\mathrm{d}^{2}\alpha^{\prime}$, then the time-evolution of the Wigner function of an arbitrary state satisfies
$\begin{split}W(\alpha,t)&=W\left(\alpha_{c}(0)+\mathrm{e}^{i\omega_{0}t}(\alpha-\alpha_{c}(t)),t=0\right)\\ &=W\left(\mathrm{e}^{i\omega_{0}t}\alpha-i\sqrt{\frac{m}{2\hbar\omega_{0}}}\int_{0}^{t}a(t^{\prime})\mathrm{e}^{i\omega_{0}t^{\prime}}\mathrm{d}t^{\prime},t=0\right).\end{split}$ (6)
Thus, the shape of the Wigner function of an arbitrary state doesn’t change during time evolution, and its time evolution is only a transformation in the phase space composited by a translation $\alpha_{c}(t)$ plus a rotation with a constant angular speed $\omega_{0}$. Besides, the evolution of an arbitrary quantum state is completely determined by its classical trajectory $x_{c}(t)$ and momentum $p_{c}(t)$ satisfying the classical equations of motion (4) in the phase space.
Specially, if the interferometer is initially prepared as a coherent state |$\ket{\beta_{0}}=\hat{D}(\beta_{0})\ket{0}$, then it evolves as
|$\ket{\psi(t)}=\mathrm{e}^{\frac{i}{\hbar}\int_{0}^{t}L[x_{c}(t^{\prime}),\dot{x}_{c}(t^{\prime})]\mathrm{d}t^{\prime}}\mathrm{e}^{-i\mathrm{Im}(\Delta\alpha_{c}\beta_{0}^{*})}\mathrm{e}^{-\frac{1}{2}\omega_{0}t}\ket{\alpha_{c}(t)+\mathrm{e}^{-i\omega_{0}t}(\beta_{0}-\alpha_{c}(0))},$ (7)
with the notation $\Delta\alpha_{c}\equiv\alpha_{c}(t)-\alpha_{c}(0)$. So every single arm remains in a coherent state during the time evolution, of which the wavefunction keeps in a Gaussian wave packet with a fixed width. Besides, the time evolution of an initial coherent state can be determined by the classical trajectory. Remarkably, there are additional phases $\mathrm{Im}(\Delta\alpha_{c}\beta_{0}^{*})$ and $-\omega_{0}t/2$ related to the details of the initial quantum state.
2.2 Time evolution of interferometer and witness in spin-space
For an interferometer, the driven forces on its different arms are spin-dependent, so the corresponding time-evolution operators are different. In particular, the spatial part of the density matrix related to the spin element |$\ket{i}\bra{j}$ | evolves as $\hat{U}_{i}(t)\hat{\rho}_{0}\hat{U}^{\dagger}_{j}(t)$ with $i,j=\pm 1$. Suppose the spin state is initially prepared as a pure state |$\ket{S_{z}(t=0)}=(\ket{1}+\ket{-1})/\sqrt{2}$, then the total density matrix of the interferometer at time $t$ is
$\hat{\rho}_{\rm tot}(t)$ $=\frac{1}{2}\sum\limits_{i,j=\pm 1}\mathrm{e}^{i(\phi_{i}(t)-\phi_{j}(t))}\ket{S_{z}=i}\bra{S_{z}=j}\otimes\hat{\rho}_{ij}(t),$ (8)
where $\text{where }\hat{\rho}_{ij}(t)$ $\equiv\hat{D}(\alpha_{i}(t))\mathrm{e}^{-\frac{i}{\hbar}\hat{H}_{0}t}\hat{D}(-\alpha_{i}(0))\hat{\rho}_{0}\hat{D}(\alpha_{j}(0))\mathrm{e}^{\frac{i}{\hbar}\hat{H}_{0}t}\hat{D}(-\alpha_{j}(t)),$
where the phase $\phi_{i,j}(t)$ only contains the spatial path integral along the classical trajectories and the spin dynamics are ignored for simplicity.
At the final time $t_{f}$ of the experiment, the observable quantity is usually constructed by a witness operator $\hat{W}$ in the spin space, so the spatial wavefunctions have to be traced out, i.e. $\hat{\rho}_{\rm spin}(t)=\mathrm{Tr}_{x}[\hat{\rho}(t)]$. If the final position $\alpha_{1}(t)$ and $\alpha_{-1}(t)$ in the classical phase space are designed the same at the final time $t_{f}$, then all the spatial density matrix elements $\hat{\rho}_{ij}(t_{f})$ are the same. As a consequence, the spin space and the spatial space are separable, i.e. $\hat{\rho}_{\rm tot}(t_{f})=\hat{\rho}_{\rm spin}(t_{f})\otimes\hat{\rho}(t_{f})$, so the spin density matrix is simply
$\hat{\rho}_{\rm spin}(t_{f})=\frac{1}{2}\begin{pmatrix}1&0&\mathrm{e}^{-i\phi_{\rm diff}}\\ 0&0&0\\ \mathrm{e}^{i\phi_{\rm diff}}&0&1\end{pmatrix},$ (9)
with the notation $\phi_{\rm diff}\equiv\phi_{-1}(t_{f})-\phi_{1}(t_{f})$. A common-used witness to measure the differential phase $\phi_{\rm diff}$ is the Ramsey interferometry [48, 4, 19]. It is constructed by measuring the population of |$\ket{S_{z}=0}$ state after a $\pi/2$-pulse operating on the state. As is derived in the B, the witness operator for Ramsey interferometry in spin-1 system is given by
$\hat{W}=\begin{pmatrix}\frac{1}{2}&0&\frac{1}{2}\\ 0&0&0\\ \frac{1}{2}&0&\frac{1}{2}\end{pmatrix}.$ (10)
Then one can obtain the observable result for the Ramsey interferometry is $\mathrm{Tr}[\hat{\rho}_{\rm spin}(t_{f})\hat{W}]=\cos^{2}(\phi_{\rm diff}/2)$, which means that the population of the eigenstate |$\ket{S_{z}=0}$ at the final time after a $\pi/2$-pulse is $\cos^{2}(\phi_{\rm diff}/2)$. For a Stern-Gerlach interferometer coupling to the local gravity by $\hat{H}_{\rm int}=mg\hat{x}$ in (1), the differential phase $\phi_{\rm diff}$ is proportional to $g$, so this device can be used as a gravimeter to measure the gravitational acceleration $g$ through the witness (10) of Ramsey interferometry [4].
3 Interferometer under Acceleration Noises
In a real experiment, there always exist some acceleration noises $\delta a(t)$ such as the vibration of the mechanical support of the system, the inertial force due to the vibration of the lab (known as the residue acceleration noise (RAN) [21, 22]) and electric forces caused by ambient electric particles [23].
In most experimental situations, acceleration noises can be described by a (wide-sense) stationary Gaussian process with a zero mean value. In particular, at any time $t$, the acceleration noise $\delta a(t)$ follows a Gaussian distribution. In addition, the wide-sense stationarity requires the auto-correlation function of $\delta a(t)$ doesn’t vary with respect to time $t$, i.e. $\mathbb{E}[\delta a(t)\delta a(t+\tau)]$ only depends on $\tau$, where $\mathbb{E}[\cdot]$ represents the statistical average of stochastic variables. According to the Wiener-Khinchin theorem, the auto-correlation function of $\delta a(t)$ is the Fourier transform of the corresponding power spectrum density (PSD)
$\mathbb{E}[\delta a(t)]$ $=0,\quad\forall t,$ (11)
$\mathbb{E}[\delta a(t)\delta a(t+\tau)]$ $=\int S_{aa}(\omega)\mathrm{e}^{-i\omega\tau}\mathrm{d}\omega,\quad\forall t,$
where $S_{aa}(\omega)=\lim\limits_{T\to\infty}|\delta a_{T}(\omega)|^{2}/T$ and $\delta a_{T}(\omega)$ is the Fourier transform of $\delta a(t)$ on the finite time domain $0\sim T$. Note that the second equation is independent with $t$ because the noise is assumed stationary. In addition, let $\tau=0$, then the auto-correlation function of $\delta a(t)$ gives the variance of $\delta a(t)$ as $\sigma_{a}^{2}(t)\equiv\mathbb{E}[(\delta a(t))^{2}]=\int S_{aa}(\omega)\mathrm{d}\omega$, which is also time-independent.
The acceleration noise $\delta a(t)$ couples to the test mass directly via the Hamiltonian $\hat{H}_{\rm noise}=-m\delta a(t)\hat{x}$, then the Hamiltonian (2) for a single arm becomes to
$\hat{H}(t)=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}-ma(t)\hat{x}-m\delta a(t)\hat{x},$ (12)
which is still of the form as (2). Hence, the time evolution operator (3) also holds when the noise is present, although the phase and the complex coordinate $\alpha_{c}(t)$ fluctuate. As a consequence, the displacement operators $\hat{D}(\alpha_{c}(0))$ keeps the same and $\hat{D}(\alpha_{c}(t))\to\hat{D}(\alpha_{c}(t)+\delta\alpha(t))$.
There are also some other noises coupling to the interferometer in different ways such as the gravity gradient noise (GGN) described by $m\delta\omega^{2}(t)\hat{x}^{2}/2$ [21, 22] and the fluctuation of the magnetic field described by $g\mu_{B}\delta\eta(t)\hat{S}_{z}\hat{x}$ with $\eta$ denoting the gradient of the magnetic field $\eta\equiv\partial B(t)/\partial x$. But these noises can be approximately regarded as acceleration noises by treating some operators as c-numbers. In particular, one may define $\delta a(x,t)=\delta\omega^{2}(t)x/2$ and $\delta a(s_{z},t)=g\mu_{B}s_{z}\delta\eta(t)/m$ with a c-number $s_{z}=\pm 1$, then the Hamiltonians for GGN and magnetic field noise become $m\delta a(x,t)\hat{x}$ and $m\delta a(s_{z},t)\hat{x}$. Thus, the theory for acceleration noises can be used to study the leading-order effects of these noises. Further analysis will be discussed in the next section.
In this section, I will discuss the aforementioned mechanisms of decoherence induced by acceleration noises. Subsection 3.1 will develop the linear response theory for dephasing. Subsection 3.2 will analyse the Humpty-Dumpty problem for a single arm, including the fluctuation of the classical trajectories and the resulting fidelity loss. Subsection 3.3 will provide a brief review on the position-localisation decoherence induced by classical acceleration noise. Finally, subsection 3.4 will apply the results on spatial degrees of freedom to the spin decoherence.
3.1 Fluctuation of path integral phase $\delta\phi$
According to the time-evolution operator (3), the spatial phase is precisely the path integral along the classical trajectories of the centre of mass. When the test mass experiences disturbances due to an acceleration noise, this phase is generally fluctuated by two leading-order channels. The first channel involves the path integral of the Lagrangian $m\delta a(t)x$ along the classical trajectory $x_{c}(t)$, while the second channel arises from the fluctuation of the classical trajectory $\delta x$ and velocity $\delta\dot{x}$ in the path integral of the ideal Lagrangian. The interplay between these two channels represents a higher-order effect and can be neglected. Thus, the phase fluctuation of a single arm can be expressed as
$\delta\phi=\frac{1}{\hbar}\int m\dot{x}\delta\dot{x}-m\omega_{0}^{2}x\delta x+ma(t)\delta x+m\delta a(t)x\mathrm{d}t.$ (13)
In order to evaluate the phase fluctuation $\delta\phi$, it is notable that the dynamical equation (4) can be written as another form as $\ddot{x}+\omega_{0}^{2}x=a(t)$, then one can find out an identity
$m\dot{x}\delta\dot{x}-m\omega_{0}^{2}x\delta x+ma(t)\delta x=m\frac{\mathrm{d}}{\mathrm{d}t}\left(\dot{x}\delta x\right).$ (14)
Thus, the phase fluctuation arising from the fluctuation of the trajectory $\delta x(t)$ and velocity $\delta\dot{x}(t)$ is a total derivative that will vanish after the path integral. Remarkably, some authors have already noticed that this channel of the phase fluctuation has to be taken into account in principle, but they neglected this effect by assuming it is a higher-order effect [21, 22, 23, 24]. As is proved above, the dynamical equation ensures this effect is strictly negligible without any assumptions  [^1] This conclusion can be generalized to a generic matter-wave interferometer under a general noise. Suppose the interferometer is built up through a Lagrangian $L[x_{i}(t),\dot{x}_{i}(t)]$ and the noise is described by a Lagrangian $\delta L[x_{i}(t),\dot{x}_{i}(t)]$. Denote the ideal trajectory as $x_{i}^{0}(t)$, and the ideal phase is $\phi^{0}=\int L[x_{i}^{0},\dot{x}_{i}^{0}]/\hbar\mathrm{d}t$. Then the phase under noise at the leading order is $\begin{split}\phi^{0}+\delta\phi&=\frac{1}{\hbar}\int L[x_{i}^{0}+\delta x_{i},\dot{x}_{i}^{0}+\delta\dot{x}_{i}^{0}]+\delta L[x_{i}^{0},\dot{x}_{i}^{0}]\mathrm{d}t\\ &=\frac{1}{\hbar}\int L[x_{i}^{0},\dot{x}_{i}^{0}]+\frac{\partial L}{\partial x_{i}}\bigg{|}_{(x_{i}^{0},\dot{x}_{i}^{0})}\delta x_{i}+\frac{\partial L}{\partial\dot{x}_{i}}\bigg{|}_{(x_{i}^{0},\dot{x}_{i}^{0})}\delta\dot{x}_{i}+\delta L[x_{i}^{0},\dot{x}_{i}^{0}]\mathrm{d}t.\end{split}$ According to the Euler-Lagrangian equation $\partial L/\partial x_{i}=\mathrm{d}(\partial L/\partial\dot{x}_{i})/\mathrm{d}t$, one can prove an identity $\frac{\partial L}{\partial x_{i}}\bigg{|}_{(x_{i}^{0},\dot{x}_{i}^{0})}\delta x_{i}+\frac{\partial L}{\partial\dot{x}_{i}}\bigg{|}_{(x_{i}^{0},\dot{x}_{i}^{0})}\delta\dot{x}_{i}=\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial L}{\partial\dot{x}_{i}}\bigg{|}_{(x_{i}^{0},\dot{x}_{i}^{0})}\delta x_{i}\right).$ Thus, the phase fluctuation arising from the fluctuation of the trajectory $\delta x_{i}$ and the velocity $\delta\dot{x}_{i}$ is the integral of a total derivative. Then one may read $\delta\phi=\int\delta L[x_{i}^{0},\dot{x}_{i}^{0}]/\hbar\mathrm{d}t$ by neglecting this total derivative term. Note that the same conclusion has been proved by using Feynman’s path integral method [49]. . Therefore, the phase fluctuation of a single path is simply
$\delta\phi(t)=\frac{1}{\hbar}\int_{0}^{t}m\delta a(t^{\prime})x(t^{\prime})\mathrm{d}t^{\prime}.$ (15)
So the phase fluctuation $\delta\phi(t)$ of a single path is an Ito integral of the stochastic noise $\delta a(t)$ and can be regarded as a linear response to $\delta a(t)$. Since the acceleration noise $\delta a(t)$ is assumed as a stationary Gaussian process, the phase fluctuation $\delta\phi(t)$ is a Gaussian process with a zero mean-value $\mathbb{E}[\delta\phi(t)]=0$ for arbitrary time $t$, although it is no longer stationary. The variance of $\delta\phi(t)$ can be computed as
$\sigma_{\phi}^{2}(t)\equiv\mathbb{E}[\left(\delta\phi(t)\right)^{2}]=\frac{m^{2}}{\hbar^{2}}\int S_{aa}(\omega)|x_{t}(\omega)|^{2}\mathrm{d}\omega,$ (16)
where $x_{t}(\omega)$ is the Fourier transform of the trajectory $x(t^{\prime})$ in the finite time domain $0\sim t$, which can be written by a convolution between $x(\omega)$ and the sinc-function as $x_{t}(\omega)=x(\omega)\ast\text{sinc}(\omega t)$. Furthermore, $x(\omega)$ plays a role of the transfer function (also known as frequency response function or filter function) of the linear response of $\delta\phi$ to $\delta a$ [50, 22, 24, 44].
For the interferometer with two paths, the phase fluctuations of both arms are linear responses to the acceleration noises, so the differential phase $\delta\phi_{\rm diff}$ is also a linear response to $\delta a$ and its variance is
$\delta\phi_{\rm diff}$ $=\frac{m}{\hbar}\int_{0}^{t_{f}}\delta a(t)(x_{1}(t)-x_{-1}(t))\mathrm{d}t,$ (17)
$\sigma_{\phi_{\rm diff}}^{2}$ $=\frac{m^{2}}{\hbar^{2}}\int S_{aa}(\omega)F(\omega)\mathrm{d}\omega,$
where $F(\omega)=|x_{1}(\omega)-x_{-1}(\omega)|^{2}$ is the transfer function or filter function [50, 22, 24, 44]. Generally speaking, $F(\omega)$ tends constant in the low-frequency limit and decreases as $\omega^{-2k}$ for some index $k$ in the high-frequency limit. In particular, in the low-frequency limit $\omega\to 0$, the factor $\mathrm{e}^{i\omega t}\to 1$, then the Fourier transform satisfies $x(\omega)=\int x(t)\mathrm{e}^{i\omega t}\mathrm{d}t\approx\int x(t)\mathrm{d}t$ which is independent with $\omega$, so $F(\omega)$ is approximately constant to $\omega$ in the low-frequency limit. In the high-frequency limit, since the trajectories can be expanded as Taylor’s series as $x_{1}(t)-x_{-1}(t)=\sum_{n=k}^{\infty}c_{n}t^{n}$ with the leading order as $t^{k}$, then the corresponding Fourier transform is $x_{1}(\omega)-x_{-1}(\omega)=\sum_{n=k}^{\infty}c_{n}\omega^{-n-1}$, so the transfer function is proportional to $\omega^{-2k-2}$ at the leading order in the high-frequency limit.
Besides, recall that the dynamical equation of the trajectories is $\ddot{x}_{i}+\omega_{0}^{2}=a_{i}(t)$, which becomes to $(-\omega^{2}+\omega_{0}^{2})x_{i}(\omega)=a_{i}(\omega)$ in frequency space, so the transfer function can be written as
$F(\omega)=\frac{|a_{1}(\omega)-a_{-1}(\omega)|^{2}}{(\omega^{2}-\omega_{0}^{2})^{2}}.$ (18)
If $a_{i}(\omega)$ has no singularity, the transfer function is a meromorphic function with two second-order poles ±$\pm\omega_{0}$. Since the $S_{aa}(\omega)$ and $a_{i}(\omega)$ are symmetric in the frequency domain, the variance of the phase fluctuation can be computed by the residue theorem as
$\sigma_{\phi_{\rm diff}}^{2}=4\pi\frac{m^{2}}{\hbar^{2}}\frac{\mathrm{d}}{\mathrm{d}\omega}\left[\frac{S_{aa}(\omega)}{(\omega+\omega_{0})^{2}}\left|a_{1}(\omega)-a_{-1}(\omega)\right|^{2}\right]\bigg{|}_{\omega=\omega_{0}}.$ (19)
Suppose the math properties of $a_{i}(\omega)$ and $S_{aa}(\omega)$ are good enough such that their derivatives don’t have dominant contributions to $\sigma_{\phi_{\rm diff}}^{2}$, then $\sigma_{\phi_{\rm diff}}^{2}$ is approximately proportional to $S_{aa}(\omega_{0})$, i.e. $\sigma_{\phi_{\rm diff}}^{2}\approx m^{2}AS_{aa}(\omega_{0})/\hbar^{2}$ with the factor $A=\pi|a_{1}(\omega_{0})-a_{-1}(\omega_{0})|^{2}/\omega_{0}^{3}$. In this sense, the transfer function $F(\omega)$ is approximately a delta function, i.e. $F(\omega)\approx A\delta(\omega-\omega_{0})$.
Figure 2: (a) The unperturbed trajectories of an interferometer under a time-independent magnetic field $B=\eta x$. (b) The corresponding transfer function for the trajectories with $N$ periods.
For example, consider a time-independent magnetic field $B(t)=\eta x$ with $\eta\equiv\partial B/\partial x$. As is shown in Fig. 2 (a), the unperturbed trajectories are $x_{\pm}(t)=A_{\pm}(1-\cos\omega_{0}t)$ with the amplitute $A_{\pm}=\pm g_{\rm NV}\mu_{B}\eta/(m\omega_{0}^{2})$, where the gravitational forces are ignored for simplicity. If the total experiment time is $2\pi N/\omega_{0}$, i.e. the trajectories repeat $N$ periods in a single run of the experiment, then the corresponding transfer function is
$F_{N}(\omega)=\frac{4(\Delta A)^{2}\omega_{0}^{4}\sin^{2}\frac{N\pi\omega}{\omega_{0}}}{\omega^{2}(\omega^{2}-\omega_{0}^{2})},$ (20)
where $\Delta A\equiv A_{+}-A_{-}$. According to (20), $F_{N}(\omega)$ tends to a constant $4\pi^{2}N^{2}(\Delta A)^{2}/\omega_{0}^{2}$ in the low-frequency limit $\omega\to 0$ and decreases with respect to $\omega^{-6}$ in the high-frequency limit $\omega\to\infty$. As is shown in Fig. 2 (b), the transfer function has a peak at $\omega=\omega_{0}$, so it is approximated as a delta function $\delta(\omega-\omega_{0})$, except when the number of periods $N=1$, in which case this peak is hidden. Therefore, the variance of the phase fluctuation in this system can be estimated as
$\sigma_{\phi_{\rm diff}}^{2}\approx 4\pi m^{2}\frac{N^{2}(\Delta A)^{2}S_{aa}(\omega_{0})}{\omega_{0}\hbar^{2}}.$ (21)
3.2 Fluctuation of classical phase space $\delta\alpha$ and fidelity loss of single arm
As is discussed at the beginning of this section, the time-evolution operator of the interferometer under the noise $\delta a(t)$ has the same form as the ideal situation. Thus, the time evolution (6) of the Wigner function of an arbitrary initial state becomes
$W(\alpha,t)=W(\alpha_{c}(0)+\mathrm{e}^{i\omega_{0}t}(\alpha-\alpha_{c}(t)-\delta\alpha(t)),t=0).$ (22)
The first notable conclusion is that the acceleration noise $\delta a(t)$ doesn’t change the shape of the Wigner function of an arbitrary initial state, and it only affects the path integral phase $\delta\phi$ and the classical trajectory $\delta\alpha$ in phase space. Specially, for a coherent state given by (7), the wavefunction under the acceleration noise is still a coherent state, although its parameter $\beta(t)=\alpha_{c}(t)+\mathrm{e}^{-i\omega_{0}t}(\beta_{0}-\alpha_{c}(0))$ changes to $\beta^{\prime}(t)=\beta(t)+\delta\alpha(t)$. According to the time-evolution (5) of the complex coordinate $\alpha_{c}$ in the phase space, its fluctuation $\delta\alpha$ under the acceleration noise $\delta a(t)$ is given by
$\delta\alpha(t)=i\sqrt{\frac{m}{2\hbar\omega_{0}}}\mathrm{e}^{i\omega_{0}t}\int_{0}^{t}\delta a(t^{\prime})\mathrm{e}^{-i\omega_{0}t^{\prime}}\mathrm{d}t^{\prime}.$ (23)
So $\delta\alpha(t)$ is also a linear response to the acceleration noise $\delta a(t)$. Besides, $\delta\alpha(t)$ is also a Gaussian process with a zero mean-value $\mathbb{E}[\delta\alpha(t)]=0$ and its auto-correlation function is
$\mathbb{E}[\delta\alpha(t)\delta\alpha^{*}(t+\tau)]=\frac{m}{2\hbar\omega_{0}}\mathrm{e}^{-\frac{i}{2}\omega_{0}\tau}t(t+\tau)\int S_{aa}(\omega)\mathrm{e}^{-\frac{i}{2}\omega\tau}\text{sinc}\frac{(\omega-\omega_{0})t}{2}\text{sinc}\frac{(\omega-\omega_{0})(t+\tau)}{2}\mathrm{d}\omega,$ (24)
Remarkably, $\mathbb{E}[\delta\alpha(t)\delta\alpha^{*}(t+\tau)]$ relies on both $t$ and $\tau$, so $\delta\alpha$ is not stationary. Let $\tau=0$, then one may have the variance of $\delta\alpha$ at any time $t$ as
$\sigma_{\alpha}^{2}(t)\equiv\mathbb{E}[|\delta\alpha(t)|^{2}]=\frac{m}{2\hbar\omega_{0}}t^{2}\int S_{aa}(\omega)\text{sinc}^{2}\frac{(\omega-\omega_{0})t}{2}\mathrm{d}\omega.$ (25)
Thus, $\delta\alpha$ is a kind of diffusion process. During the beginning of the diffusion process  [^2] Here, asymptotic behaviours of the sinc-function have been used. That is, $\text{sinc}^{2}((\omega-\omega_{0})t/2)=1+\mathcal{O}(t^{2})$ for small $t$, and $\text{sinc}^{2}((\omega-\omega_{0})t/2)=2\pi\delta(\omega-\omega_{0})/t$ for large $t$., $\delta\alpha$ diffuses as $\sigma_{\alpha}^{2}(t)=D_{1}t^{2}$ with a generalized diffusion parameter $D_{1}=m\sigma_{a}^{2}/(2\hbar\omega_{0})$. In the long time limit, $\delta\alpha$ diffuses as $\sigma_{\alpha}^{2}(t)=D_{2}t$, which is similar to the standard Brownian motion with a diffusion parameter $D_{2}=\pi mS_{aa}(\omega_{0})/(\hbar\omega_{0})$. In other words, the diffusion process of $\delta\alpha$ in the long time limit is Markovian or memoryless, which means its long-time diffusion is independent of the history of the beginning diffusion.
In frequency space, the dynamical equation (4) of $\alpha_{c}$ implies a frequency response relation as $\delta\alpha(\omega)=\chi(\omega)\delta a(\omega)$, where the factor $\chi(\omega)=\sqrt{m/(2\hbar\omega_{0})}/(\omega-\omega_{0})$ is a generalized mechanical susceptibility of $\alpha_{c}(\omega)$. Then the PSD of $\delta\alpha$ is
$S_{\alpha\alpha}(\omega)=\frac{m}{2\hbar\omega_{0}}\frac{S_{aa}(\omega)}{(\omega-\omega_{0})^{2}}.$ (26)
It is notable that the Wiener-Khinchin theorem doesn’t hold for $\delta\alpha$ because it is not stationary, and one may check that the relationship between the auto-correlation function and PSD of $\delta\alpha$ is
$\mathbb{E}[\delta\alpha(t)\delta\alpha^{*}(t+\tau)]=\int S_{\alpha\alpha}(\omega)\mathrm{e}^{-i\omega\tau}\left(1+\mathrm{e}^{-i(\omega-\omega_{0})\tau}-\mathrm{e}^{i(\omega-\omega_{0})t}-\mathrm{e}^{-i(\omega-\omega_{0})(t+\tau)}\right)\mathrm{d}\omega.$ (27)
Since the trajectory of a single arm is disturbed by the noise, the corresponding quantum state is also affected, which is usually characterised by the Uhlmann fidelity [51, 52]. In particular, denote the ideal and noisy density matrices of the single arm at an arbitrary time $t$ as $\hat{\rho}_{\rm id}(t)\equiv\hat{U}_{\rm id}(t)\hat{\rho}_{0}\hat{U}_{\rm id}^{\dagger}(t)$ and $\hat{\rho}_{\rm n}(t)\equiv\hat{U}_{\rm n}(t)\hat{\rho}_{0}\hat{U}_{\rm n}^{\dagger}(t)$, where $\hat{U}_{\rm id}(t)$ and $\hat{U}_{\rm n}(t)$ are given by (3) with complex coordinates $\alpha_{c}(t)$ and $\alpha_{c}(t)+\delta\alpha(t)$, then the corresponding fidelity is defined as
$F(\hat{\rho}_{\rm id}(t),\hat{\rho}_{\rm n}(t))\equiv\text{Tr}\left[\sqrt{\sqrt{\hat{\rho}_{\rm id}(t)}\hat{\rho}_{\rm n}(t)\sqrt{\hat{\rho}_{\rm id}(t)}}\right].$ (28)
Notice that $\sqrt{\hat{\rho}_{\rm id}(t)}=\hat{U}_{\rm id}(t)\sqrt{\hat{\rho}_{0}}\hat{U}_{\rm id}^{\dagger}(t)$ and $\hat{U}_{\rm id}^{\dagger}(t)\hat{U}_{\rm n}(t)=\hat{D}\left(\delta\alpha(t)\mathrm{e}^{i\omega_{0}t}\right)\mathrm{e}^{i\delta\phi(t)}$ with the phase $\delta\phi(t)$ including the path integral phase and quantum phase arising from the commutation between operators, then it can be proved that
$F(\hat{\rho}_{\rm id}(t),\hat{\rho}_{\rm n}(t))=F\left(\hat{\rho}_{0},\hat{D}(\delta\alpha(t)\mathrm{e}^{i\omega_{0}t})\hat{\rho}_{0}\hat{D}(-\delta\alpha(t)\mathrm{e}^{i\omega_{0}t})\right).$ (29)
So the fidelity between $\hat{\rho}_{\rm id}(t)$ and $\hat{\rho}_{\rm n}(t)$ is exactly the fidelity of the initial state $\hat{\rho}_{0}$ and its fluctuation under $\hat{D}(\delta\alpha(t)\mathrm{e}^{i\omega_{0}t})$. Specially, if the system is initially prepared as a pure state $\hat{\rho}_{0}=\ket{\psi_{0}}\bra{\psi_{0}}$ |, then the fidelity is simply the contrast
$F(\hat{\rho}_{\rm id}(t),\hat{\rho}_{\rm n}(t))=|\bra{\psi_{0}}\hat{D}(\delta\alpha(t)\mathrm{e}^{i\omega_{0}t})\ket{\psi_{0}}|.$ | . (30)
In addition, if the system is initially prepared as a coherent state, then the fidelity is
$F(\hat{\rho}_{\rm id}(t),\hat{\rho}_{\rm n}(t))=\exp\left[-\frac{|\delta\alpha(t)|^{2}}{2}\right].$ (31)
3.3 Position localisation decoherence
It is notable that the randomness of the trajectory fluctuation $\delta\alpha(t)$ will cause a position localisation decoherence [10] on the single arm of the interferometer [12, 13]. For an ideal interferometer, the matrix element of the density matrix $\hat{\rho}_{\rm id}(t)=\hat{U}_{\rm id}(t)\hat{\rho}_{0}\hat{U}_{\rm id}^{\dagger}(t)$ satisfies
$\bra{x_{1}}\hat{\rho}_{\rm id}(t)\ket{x_{2}}=\bra{x^{\prime}_{1}(t)}\hat{\rho}_{0}\ket{x^{\prime}_{2}(t)},$ (32)
$x^{\prime}_{i}(t)=\left(x_{i}-\sqrt{\frac{2\hbar}{m\omega_{0}}}\text{Re}\alpha(t)\right)\cos\omega_{0}t+\sqrt{\frac{2\hbar}{m\omega_{0}}}\text{Re}\alpha(0).$
So the time-evolution of the density matrix element is simply the oscillation defined by the Hamiltonian (2) without noise, which has been concluded by the Wigner function (6). Besides, one can prove that the purity of the density matrix $\mathcal{P}_{\rm id}(t)\equiv\text{Tr}\left[\hat{\rho}_{\rm id}^{2}(t)\right]$ keeps invariant according to the following calculation
$\begin{split}\mathcal{P}_{\rm id}(t)&=\int\mathrm{d}x_{1}\mathrm{d}x_{2}\bra{x_{1}}\hat{\rho}_{\rm id}(t)\ket{x_{2}}\bra{x_{2}}\hat{\rho}_{\rm id}(t)\ket{x_{1}}\\ &=\int\mathrm{d}x^{\prime}_{1}(t)\mathrm{d}x^{\prime}_{2}(t)\bra{x^{\prime}_{1}(t)}\hat{\rho}_{0}\ket{x^{\prime}_{2}(t)}\bra{x^{\prime}_{2}(t)}\hat{\rho}_{0}\ket{x^{\prime}_{1}(t)}=\text{Tr}\left[\hat{\rho}_{0}^{2}\right]\equiv\mathcal{P}_{0}.\end{split}$ (33)
On the other hand, for the interferometer under noises, according to the identity $\hat{U}_{\rm n}(t)=\hat{D}(\delta\alpha(t))\hat{U}_{\rm id}(t)\mathrm{e}^{i\delta\phi(t)}$ with some phase $\delta\phi(t)$, the matrix element of the density matrix $\hat{\rho}_{\rm n}(t)=\hat{U}_{\rm n}(t)\hat{\rho}_{0}\hat{U}_{\rm n}^{\dagger}(t)$ is
$\bra{x_{1}}\hat{\rho}_{\rm n}(t)\ket{x_{2}}=\bra{x^{\prime}_{1}(t)}\hat{D}(\delta\alpha(t))\hat{\rho}_{0}\hat{D}^{\dagger}(\delta\alpha(t))\ket{x^{\prime}_{2}(t)}.$ (34)
To compute the ensemble average of this matrix element, one can consider the Glauber’s P-representation [47] of the initial density matrix $\hat{\rho}_{0}=\int P(\beta)\ket{\beta}\bra{\beta}\mathrm{d}^{2}\beta$. According to the identity $\hat{D}(\delta\alpha(t))\ket{\beta}=\mathrm{e}^{i\phi(\beta,\delta\alpha)}\ket{\beta+\delta\alpha(t)}$, the matrix element of $\hat{\rho}_{\rm n}(t)$ becomes to
$\bra{x_{1}}\hat{\rho}_{\rm n}(t)\ket{x_{2}}=\int P(\beta)\braket{x^{\prime}_{1}(t)}{\beta+\delta\alpha(t)}\braket{\beta+\delta\alpha(t)}{x^{\prime}_{2}(t)}\mathrm{d}^{2}\beta.$ (35)
Based on the wavefunction of a coherent state, one can obtain the ensemble average of the matrix element in the P-representation satisfying $\mathbb{E}\left[\braket{x^{\prime}_{1}(t)}{\beta+\delta\alpha(t)}\braket{\beta+\delta\alpha(t)}{x^{\prime}_{2}(t)}\right]=\braket{x^{\prime}_{1}(t)}{\beta}\braket{\beta}{x^{\prime}_{2}(t)}\mathrm{e}^{-\frac{m\omega_{0}}{\hbar}\sigma_{\alpha}^{2}(t)(x^{\prime}_{1}(t)-x^{\prime}_{2}(t))^{2}}$. Therefore, the density matrix decoheres as
$\bra{x_{1}}\mathbb{E}\left[\hat{\rho}_{\rm n}(t)\right]\ket{x_{2}}=\bra{x^{\prime}_{1}(t)}\hat{\rho}_{0}\ket{x^{\prime}_{2}(t)}\mathrm{e}^{-\frac{m\omega_{0}}{\hbar}\sigma_{\alpha}^{2}(t)(x^{\prime}_{1}(t)-x^{\prime}_{2}(t))^{2}}.$ (36)
Consequently, in the long-time limit, off-diagonal terms of $\hat{\rho}_{\rm n}(t)$ decays exponentially and only the diagonal terms of $\hat{\rho}_{\rm n}(t)$ keeps non-zero, so the corresponding purity approaches to zero. Furthermore, since the variance of $\delta\alpha$ is approximately proportional to $t$ in the long time limit, then there is a well-defined decoherence rate $\gamma=\pi m^{2}(x_{1}-x_{2})^{2}S_{aa}(\omega_{0})/\hbar^{2}$. It is noteworthy that a decoherence rate proportional to $(x_{1}-x_{2})^{2}$ is related to the following master equation [10, 12]
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{\rho}=\frac{1}{i\hbar}[\hat{H},\hat{\rho}]-\Lambda[\hat{x},[\hat{x},\hat{\rho}]],$ (37)
where $\Lambda=\pi m^{2}S_{aa}(\omega_{0})/\hbar^{2}$, which is the same as [12] up to a factor $2\pi$. Such master equation can be derived directly from the Hamiltonian (2), and the math details are summarized in C. It is noteworthy that the decoherence due to the environmental scattering in the long-wavelength limit also follows the same master equation [25, 10], so the scattering of long-wavelength particles like photons can be regarded as a classical stochastic noise in a long-time limit.
3.4 Loss of witness
Figure 3: The trajectories affected by acceleration noises, where (a) for a single path, (b) for acceleration noises and (c) for path-dependent noises. As is shown in (b), under common mode noises, the final points in phase space of both arms translate the same, so the two paths can be overlapped properly. As is shown in (c), under path-dependent noises, the points of both arms translate differently, so there is a Loschmidt echo loss.
For an interferometer with two arms, the trajectory fluctuations $\delta\alpha(t)$ of both paths are the same according to (23) as long as both arms couples to the noise in the same way. As is proved by (22), the Wigner functions of the arms are totally determined by their classical trajectories in phase space, so the two arms can meet with each other properly at the final time, that is, the contrast of the two arms keeps 1, although the fidelity of each arm gets lost, shown as Fig.3 (b). In addition, the position localisation decoherence also has no influence on decoherence in spin space, because $\hat{\rho}_{\rm spin}$ only relies on the diagonal terms of $\hat{\rho}_{\rm n}(t)$ which don’t decay during evolution according to (36).
Thus, the spin space and the coordinate space are still separable, and the noise only contributes a fluctuation of the differential phase $\delta\phi_{\rm diff}$ to the ideal density matrix (9) in the spin space. It can be regarded as a type of common mode noise cancellation, which has been applied in atomic interferometers (AI) [53, 54, 55]. Notably, this effect is usually used to reduce dephasing in AIs, while it mitigates position localisation decoherence and fidelity loss in SGIs.
As discussed in sub-section 3.1, the phase fluctuations of both arms follow Gaussian distributions, so the ensemble average of the random phase factor contributes an exponential decay factor $\mathbb{E}[\mathrm{e}^{i\delta\phi_{\rm diff}}]=\mathrm{e}^{-\sigma_{\rm diff}^{2}/2}$. Then the density matrix in spin space under the ensemble average is
$\mathbb{E}[\hat{\rho}_{\rm spin}(t_{f})]=\frac{1}{2}\begin{pmatrix}1&0&\mathrm{e}^{-\frac{\sigma_{\phi_{\rm diff}}^{2}}{2}}\mathrm{e}^{-i\phi_{\rm diff}}\\ 0&0&0\\ \mathrm{e}^{-\frac{\sigma_{\phi_{\rm diff}}^{2}}{2}}\mathrm{e}^{i\phi_{\rm diff}}&0&1\end{pmatrix}.$ (38)
Recall that ideal value of the witness of the Ramsey interferometry is $\mathrm{Tr}[\hat{\rho}_{\rm spin}(t_{f})\hat{W}]=\cos^{2}(\phi_{\rm diff}/2)$. For the spin density matrix (38) under noise, the observable result of the witness and the corresponding loss are
$\braket{W}$ $\equiv\mathrm{Tr}\left[\mathbb{E}[\hat{\rho}_{\rm spin}(t_{f})]\hat{W}\right]=\frac{1}{2}\left[1+\mathrm{e}^{-\frac{\sigma_{\phi_{\rm diff}}^{2}}{2}}\cos\phi_{\rm diff}\right],$ (39)
$\Delta W$ $=\frac{\mathrm{e}^{-\frac{\sigma_{\phi_{\rm diff}}^{2}}{2}}-1}{2}\cos\phi_{\rm diff}\approx-\frac{\sigma_{\phi_{\rm diff}}^{2}}{2}\cos\phi_{\rm diff}.$
Therefore, the loss of witness is proportional to the variance of the phase fluctuation. Recall that $\sigma^{2}_{\phi_{\rm diff}}$ is approximately proportional to $S_{aa}(\omega_{0})$, so the loss of witness $\Delta W$ should be mitigated by controlling the PSD at the intrinsic frequency $\omega_{0}$ of the interferometer in a real experiment.
Drawing from the discussion in this section, I conducted a theoretical analysis of the influence of acceleration noise on an interferometer. The spatial decoherence encompasses dephasing, fidelity loss and position-localisation decoherence, with the key points outlined as follows:
  • Noises can induce fluctuations of the phase of each arm as a linear response, and the transfer function (or the filter function) is exactly the Fourier transform of the unperturbed classical trajectory.
  • Noises can also affect the classical trajectories in the phase space as a generalized diffusion process which approaches to a standard Brownian motion in the long-time limit, leading to a loss of fidelity of each arm, even though the shape of the Wigner function of each arm of the interferometer keeps invariant before performing the ensemble average of the noise.
  • The ensemble average of the noise can result in position localisation decoherence on each arm, characterised by a dissipator $[\hat{x},[\hat{x},\hat{\rho}]]$, resulting in a decay factor $\mathrm{e}^{-\Lambda(x_{1}-x_{2})^{2}t}$ to the spatial density matrix, with the decoherence rate $\Lambda$ is proportional to the PSD $S_{aa}(\omega_{0})$.
However, it is important to note that neither the fidelity loss nor the position localisation decoherence causes decoherence in the spin space for common mode noise. As a result, the spin decoherence only relies on the dephasing effect, and the loss of the witness is approximately proportional to $\sigma^{2}_{\phi_{\rm diff}}$.
4 Discussion and Conclusion
4.1 Path-dependent noises
Apart from acceleration noises, there exist noises coupling to the interferometer in other ways such as $m\delta\omega^{2}(t)\hat{x}^{2}$ or $\mu\delta\eta(t)\hat{S}_{z}\hat{x}$ with $\eta(t)\equiv\partial B(t)/\partial x$. In order to study these noises analytically, it is necessary in principle to construct the corresponding Lewis-Riesenfeld invariant for these coupling Hamiltonians, which can be a great challenge in math. However, by treating some operators as c-numbers, these noises can be regarded as path-dependent acceleration noises. Then the effects of these noises can be studied through the analytic solution for acceleration noises, although some higher-order properties can be lost.
As is shown in Fig.3 (c), the final states of the two paths are usually not able to meet with each other, i.e. $\alpha_{1}(t_{f})\neq\alpha_{-1}(t_{f})$, so the total density matrix is not separable. As a consequence, apart from the dephasing effect characterised by a decay factor $\mathrm{e}^{-\sigma_{\phi_{\rm diff}}^{2}/2}$, the off-diagonal terms of the spin density matrix have an additional decay factor arising from the partial trace of the spatial degrees of freedom. Therefore, the spin density matrix under path-dependent noises is given by
$\mathbb{E}[\hat{\rho}_{\rm spin}(t_{f})]=\frac{1}{2}\begin{pmatrix}1&0&\mathbb{E}\left[\text{Tr}[\hat{\rho}_{1,-1}(t_{f})]\right]\mathrm{e}^{-\frac{\sigma_{\phi_{\rm diff}}^{2}}{2}}\mathrm{e}^{-i\phi_{\rm diff}}\\ 0&0&0\\ \mathbb{E}\left[\text{Tr}[\hat{\rho}_{-1,1}(t_{f})]\right]\mathrm{e}^{-\frac{\sigma_{\phi_{\rm diff}}^{2}}{2}}\mathrm{e}^{i\phi_{\rm diff}}&0&1\end{pmatrix}.$ (40)
For path-dependent noises, the fluctuation of the differential phase of the interferometer is given by
$\delta\phi_{\rm diff}=\frac{m}{\hbar}\int_{0}^{t_{f}}(\delta a_{1}(t)x_{1}(t)-\delta a_{-1}(t)x_{-1}(t))\mathrm{d}t.$ (41)
In this case, the variance of the phase fluctuation is determined by the correlation between $\delta a_{1}(t)$ and $\delta a_{-1}(t)$. For example, if $\delta a_{1}(t)$ and $\delta a_{-1}(t)$ arise from a noise term $\delta L=mf(t)x^{n}$ with a Gaussian noise $f(t)$ by defining $\delta a_{\pm 1}(t)\equiv f(t)x_{\pm 1}^{n-1}(t)$, then the dephasing of the single arm becomes $\delta\phi_{\rm diff}=\frac{m}{\hbar}\int_{0}^{t_{f}}f(t)x^{n}(t)\mathrm{d}t$. Thus, the variance of the differential phase also has a form (17) with a transfer function $F(\omega)=|x_{1}^{n}(\omega)-x_{-1}^{n}(\omega)|^{2}$, of which the math properties are similar.
On the other hand, for a more general case, $\delta a_{1}(t)$ and $\delta a_{-1}(t)$ can be treated as independent, i.e. the cross-correlation can be approximately assumed as $\mathbb{E}[\delta a_{1}(t_{1})\delta a_{-1}(t_{2})]=0$. Then $\delta\phi_{1}$ and $\delta\phi_{-1}$ are also independent, so the variance of dephasing is
$\sigma_{\phi_{\rm diff}}^{2}=\sigma_{\phi_{1}}^{2}+\sigma_{\phi_{-1}}^{2}=\frac{m^{2}}{\hbar^{2}}\int_{0}^{t_{f}}S_{a_{1}a_{1}}(\omega)|x_{1}(\omega)|^{2}+S_{a_{-1}a_{-1}}(\omega)|x_{-1}(\omega)|^{2}\mathrm{d}\omega.$ (42)
As for the Humpty-Dumpty problem, one can compute that
$\text{Tr}[\hat{\rho}_{1,-1}(t_{f})]=\text{Tr}[\hat{D}(\Delta\alpha(t_{f})\mathrm{e}^{i\omega_{0}t_{f}})\hat{\rho}_{0}]\mathrm{e}^{i\text{Im}(\alpha_{1}(t_{f})\alpha^{*}_{-1}(t_{f}))},$ (43)
where $\Delta\alpha(t_{f})\equiv\alpha_{1}(t_{f})-\alpha_{-1}(t_{f})$. Specially, if the initial state is prepared as a Gaussian state, then the Loschmidt echo [29, 56, 57] (also referred as the contrast) is
$C\equiv\text{Tr}[\hat{D}(\Delta\alpha(t_{f})\mathrm{e}^{i\omega_{0}t_{f}})\hat{\rho}_{0}]=\mathrm{e}^{-\frac{|\Delta\alpha(t_{f})|^{2}}{2}}=\mathrm{e}^{-\frac{(\Delta x)^{2}}{m\omega_{0}/\hbar}-\frac{(\Delta p)^{2}}{m\hbar\omega_{0}}}.$ (44)
Since both $\delta\alpha_{\pm 1}$ follow Gaussian distribution, then $\delta\alpha_{1}-\delta\alpha_{-1}$ also follow Gaussian distribution with a variance $\sigma^{2}_{\rm tot}=\sigma_{1}^{2}+\sigma_{-1}^{2}$ if they are assumed independent. Note that $\sigma_{\pm 1}$ are exactly the same, then the total variance is simply $\sigma^{2}_{\rm tot}=2\sigma_{\alpha}^{2}$, where $\sigma_{\alpha}^{2}$ is exactly the variance of a single arm given by (25) which is a linear response to the acceleration noise $\delta a(t)$ with a transfer function $\text{sinc}^{2}(\omega-\omega_{0})t/2$. Thus, the ensemble average of the contrast is
$-\mathbb{E}[\log C]=\frac{1}{2}\mathbb{E}\left[|\delta\alpha_{1}-\delta\alpha_{-1}|^{2}\right]=\sigma_{\alpha}^{2}.$ (45)
To sum up, path-dependent acceleration noises can cause both dephasing and contrast loss on the interferometer, and both effects are linear responses to $\delta a(t)$, while the Humpty-Dumpty problem can lead to an additional phase $\mathrm{Im}(\alpha_{-1}^{*}(t_{f})\alpha_{1}(t_{f}))$. Both effects can contribute a decay factor on off-diagonal terms of the spin space, so the observable result for the Ramsey interferometry becomes
$\mathrm{Tr}\left[\mathbb{E}[\hat{\rho}_{\rm spin}(t_{f})]\hat{W}\right]=\frac{1}{2}\left[1+\mathrm{e}^{-\frac{\sigma_{\phi_{\rm diff}}^{2}}{2}}\mathrm{e}^{-\sigma_{\alpha}^{2}}\cos(\phi_{\rm diff}+\mathrm{Im}(\alpha_{-1}^{*}(t_{f})\alpha_{1}(t_{f})))\right].$ (46)
A final remark is that by treating some operators as c-numbers, some higher-order effects were not studied. For example, the noise described by $m\delta\omega^{2}(t)\hat{x}^{2}$ like GGN or rotation noise [24] can disorder the shape of the Wigner function by the squeezed operator $\hat{S}(\zeta)=\mathrm{e}^{(\zeta^{*}\hat{a}^{2}-\zeta\hat{a}^{\dagger 2})}$ with $\zeta(t)$ related to $\delta\omega(t)$, then the overlap of the two arms at the final time can be lost even if the centre of mass can meet with each other. As for the fluctuation of the magnetic field described by $\mu\delta\eta(t)\hat{S}_{z}\hat{x}$, this noise will affect both the spin space and the phase space, but the decoherence in the spin space cannot be studied by treating $\hat{S}_{z}$ as c-number. Therefore, it is necessary to construct the Lewis-Riesenfeld invariant for these noises to systematically study these higher-order effects in future research.
4.2 Role of randomness
It is remarkable that the density matrix before and after the ensemble average can be quite different. In particular, the phase fluctuation $\delta\phi_{\rm diff}$ contributes a pure phase factor $\mathrm{e}^{i\delta\phi_{\rm diff}}$ without ensemble average but contributes a decay factor $\mathrm{e}^{-\sigma_{\phi_{\rm diff}}^{2}}$ after average. Similarly, the contrast loss with and without the ensemble average can be also different.
To understand the role of randomness better, one can compute the purity $\text{Tr}[\hat{\rho}^{2}]$ or the von Neumann entropy $S=-\text{Tr}\left[\hat{\rho}\ln\hat{\rho}\right]$. The purities with and without the ensemble average are  [^3] The trick to compute the purity and the von Neumann entropy is to diagonalize $\hat{\rho}_{\rm spin}$. In particular, suppose the non-zero eigenvalues of $\hat{\rho}_{\rm spin}$ are $\{\lambda_{i}\}$, then the purity and the von Neumann entropy of $\hat{\rho}_{\rm spin}$ are $\sum_{i}\lambda_{i}^{2}$ and $-\sum_{i}\lambda_{i}\ln\lambda_{i}$. For the density matrix $\hat{\rho}_{\rm spin}$ with a decay factor $\mathrm{e}^{-\Gamma_{\rm tot}}$ on the off-diagonal terms, its eigenvalues are $\lambda_{1,2}=(1\pm\mathrm{e}^{-\Gamma_{\rm tot}})/2$ and $\lambda_{3}=0$. Specially, when the total decoherence factor $\Gamma_{\rm tot}$ is very small, then the entropy can be approximated as $\begin{split}S&=-\left[\frac{1+\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\ln\frac{1+\mathrm{e}^{-\Gamma_{\rm tot}}}{2}+\frac{1-\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\ln\frac{1-\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\right]\approx\frac{1+\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\cdot\frac{1-\mathrm{e}^{-\Gamma_{\rm tot}}}{2}+\frac{1-\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\cdot\frac{1+\mathrm{e}^{-\Gamma_{\rm tot}}}{2}\\ &=\frac{1}{2}(1-\mathrm{e}^{-2\Gamma_{\rm tot}})\approx\Gamma_{\rm tot},\end{split}$ where the first approximated equation used an approximation $\ln((1\pm\mathrm{e}^{-\Gamma_{\rm tot}})/2)\approx(-1\pm\mathrm{e}^{\Gamma_{\rm tot}})/2$.
$\text{Tr}[\hat{\rho}_{\rm spin}^{2}(t_{f})]$ $=\frac{1}{2}(1+C^{2})\approx 1-(-\log C),$ (47)
$\text{Tr}\left[\mathbb{E}[\hat{\rho}_{\rm spin}(t_{f})]^{2}\right]$ $=\frac{1}{2}(1+\mathbb{E}[C]^{2}\mathrm{e}^{-\sigma_{\phi_{\rm diff}}^{2}})\approx 1-\frac{\sigma_{\phi_{\rm diff}}^{2}}{2}-\sigma_{\alpha}^{2},$
and the von Neumann entropies with and without the ensemble average are
$S[\hat{\rho}_{\rm spin}]$ $\approx-\log C,$ (48)
$S[\mathbb{E}[\hat{\rho}_{\rm spin}]]$ $\approx\frac{\sigma_{\phi_{\rm diff}}^{2}}{2}+\sigma_{\alpha}^{2}.$
As is shown from the purity and the entropy, there is only the contrast loss without the ensemble average. When considering the ensemble average, the dephasing effect can cause a purity loss and an entropy increase, and the contrast loss has a different influence on purity and entropy, which can be regarded as a consequence of the position localisation decoherence.
Therefore, from the point of view of information, the randomness of the noise plays an essential role in taking out the information of the quantum system, except for the contrast loss mechanism. The information of the quantum system gets lost even if the environment is classical as long as the environment is stochastic. Thus, the information loss indicates a deep connection between the classical randomness and the quantumness [58, 59].
In summary, the information of the interferometer is lost to the classical noise through three mechanisms. The first mechanism is the dephasing, and the phase information of the two arms is lost due to the randomness of the noise. The second mechanism is the Humpty-Dumpty problem, and the information of the spin space is encoded to the spatial degrees of freedom and gets lost due to the path-dependent noises. The third mechanism is the position localisation decoherence, both randomness and path-dependence of the noise are necessary for this mechanism.
4.3 Summary
In this paper, I presented a comprehensive theoretical discussion on the decoherence of a Stern-Gerlach interferometer induced by classical stochastic acceleration noise. I first solved the time-dependent Schrodinger equation analytically using Lewis-Riesenfeld invariant and the quantum Arnold transformation to obtain the spatial motion of the test mass. Subsequently, I analysed the decoherence effects in the spin space including dephasing, position localisation decoherence and the contrast loss.
As is demonstrated in section 3, only dephasing affects the witness in spin space, which is a linear response to the noise, with the transfer function as the Fourier transform of the unperturbed trajectories. Nonetheless, the ensemble average of the noise can leads to position localisation decoherence which contributes a decay factor $\mathrm{e}^{-\Lambda(x_{1}-x_{2})^{2}t}$. Besides, the noise can also disturb the classical trajectory and velocity of every path, resulting in the fidelity loss of each individual arm. As is shown in section 3.4, position localisation decoherence and fluctuation of trajectories induced by acceleration noise don’t affect the spin witness, which can be regarded as a type of the common mode noise cancellation.
However, for more generic noises approximately regarded as path-dependent acceleration noises, the final positions and velocities of the two arms are different, which leads to a contrast loss of the quantum states and finally results in an additional decoherence in the spin space, known as the Humpty-Dumpty problem. In addition, the statistical average of the overlap of the two arms can change the value of the contrast loss, which can be regarded as a consequence of the position localisation decoherence.
Finally, it is remarkable that the randomness of the noise plays an essential role in the decoherence. Except for the Humpty-Dumpty problem, the other two mechanisms of the decoherence of the spin space, the dephasing and the position localisation decoherence, arise from the ensemble average, which has been illustrated by the purity loss and the von Neumann entropy increase. From the point of view of information theory, the information of the quantum system gets lost due to the randomness of noises through these two mechanisms. Therefore, if the acceleration noises in experiments are monitored such that the randomness is eliminated as much as possible, then the decoherence can be suppressed a lot, and the fluctuations of the phases and the trajectories become reconstructable in principle.
Appendix A Analytical Solution for the Schrodinger Equation with Acceleration Noises
In this appendix, I will show the math details to solve the Schrodinger equation with the time-dependent Hamiltonian (2) by the Lewis-Riesenfeld invariant and quantum Arnold transformation method [37, 38, 60, 40, 41].
The basic idea [39] is to find a so-called Lewis-Riesenfeld operator $\hat{I}(t)$ such that all the eigenvalues $\{\lambda_{n}\}$ of $\hat{I}(t)$ are time-independent and all the instantaneous eigenstates {|$\{\ket{n(t)}\}$ of $\hat{I}(t)$ are solutions of Schrodinger equation. Then the time-evolution operator is given by $\hat{U}(t)=\sum_{n}\mathrm{e}^{i\alpha_{n}(t)}\ket{n(t)}\bra{n(0)}$ |, where $\alpha_{n}(t)=\int_{0}^{t}\bra{n(t^{\prime})}\left(i\mathrm{d}/\mathrm{d}t^{\prime}-\hat{H}(t^{\prime})/\hbar\right)\ket{n(t^{\prime})}\mathrm{d}t^{\prime}$ is known as the Lewis-Riesenfeld phase [37, 39]. From the point view of differential geometry, the instantaneous eigenstates {|$\{\ket{n(t)}\}$ form a moving frame of the Hilbert space as a manifold, and the Lewis-Riesenfeld phase is a kind of geometric phase arising from the connection 1-form of the manifold.
Such time-dependent operator $\hat{I}(t)$ can be generally constructed by a reference frame transformation on the Hamiltonian from the lab reference to the co-moving frame [40], known as the quantum Arnold transformation [41]. In particular, for a driven oscillator with a Hamiltonian (2), consider a translation operator in the phase space of the oscillator $\hat{\mathcal{U}}(t)=\exp\left[-\frac{i}{\hbar}(\hat{x}p(t)-\hat{p}x(t))\right]=\hat{D}(-\alpha(t))$ with $\alpha(t)=\sqrt{m\omega_{0}/2\hbar}(x(t)+ip(t)/m\omega_{0})$. Under the unitary transformation, the wavefunction changes as |$\ket{\psi}\to\ket{\psi^{\prime}}=\hat{\mathcal{U}}(t)\ket{\psi}$, while the Hamiltonian changes as
$\hat{H}(t)\to\hat{\mathcal{U}}\hat{H}\hat{\mathcal{U}}^{\dagger}-i\hbar\hat{\mathcal{U}}\frac{\mathrm{d}}{\mathrm{d}t}\hat{\mathcal{U}}^{\dagger}.$ (49)
The term $-i\hbar\hat{\mathcal{U}}(\mathrm{d}\hat{\mathcal{U}}^{\dagger}/\mathrm{d}t)$, often referred as the counter-diabatic term $\hat{H}_{CD}$ [61], mathematically arises from the connection 1-form to keep the Schrodinger equation covariant under the transformation $\hat{\mathcal{U}}(t)$. If the trajectory $x(t)$ and momentum $p(t)$ is chosen satisfying the canonical equation defined by the Hamiltonian (2)
[…]
which is the real-valued form for (4), then the Hamiltonian in the co-moving reference is
$\hat{\mathcal{U}}\hat{H}\hat{\mathcal{U}}^{\dagger}-i\hbar\hat{\mathcal{U}}\frac{\mathrm{d}}{\mathrm{d}t}\hat{\mathcal{U}}^{\dagger}=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}+\delta(t),$ (51)
where the phase $\delta(t)$ is given by
$\delta(t)=-\frac{1}{2}ma(t)x_{c}(t).$ (52)
Therefore, the Hamiltonian in the comoving reference frame is time-independent except for a pure phase term $\delta(t)$. So the time-evolution of the wave function under the comoving reference is simply
|$\ket{\psi^{\prime}(t)}=\mathrm{e}^{-\frac{i}{\hbar}\int_{0}^{t}\delta(t^{\prime})\mathrm{d}t^{\prime}}\mathrm{e}^{-\frac{i}{\hbar}\left(\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}\right)t}\ket{\psi^{\prime}(0)}.$ (53)
Then the wavefunction in the original reference frame is
|$\ket{\psi(t)}=\mathrm{e}^{-\frac{i}{\hbar}\int_{0}^{t}\delta(t^{\prime})\mathrm{d}t^{\prime}}\hat{\mathcal{U}}^{\dagger}(t)\mathrm{e}^{-\frac{i}{\hbar}\left(\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\omega_{0}^{2}\hat{x}^{2}\right)t}\hat{\mathcal{U}}(0)\ket{\psi(0)}.$ (54)
Furthermore, the equation of motion (50) allows to rewrite the phase $\delta(t)$ in another form
$\delta(t)=-\left(\frac{1}{2}m\dot{x}_{c}^{2}(t)-\frac{1}{2}m\omega_{0}^{2}x_{c}^{2}(t)+ma(t)x_{c}(t)-\frac{1}{2}m\frac{\mathrm{d}}{\mathrm{d}t}(\dot{x}_{c}x_{c})\right).$ (55)
This means that the $\delta(t)$ phase term is exactly the minus Lagrangian $L[x_{c}(t),\dot{x}_{c}(t)]$ of the system (2) along the classical trajectory up to a total derivative term. Thus, the global phase in (54) is exactly the path integral along the classical trajectory $\phi=\int_{0}^{t}L[x_{c}(t^{\prime}),\dot{x}_{c}(t^{\prime})]/\hbar\mathrm{d}t^{\prime}$, so (54) presents the time evolution operator (3). Remarkably, since the $\delta(t)$ phase arises from the gauge term $-i\hbar\hat{\mathcal{U}}(\mathrm{d}\hat{\mathcal{U}}^{\dagger}/\mathrm{d}t)$, the Feynman’s path integral phase in this system has a physical interpretation of geometric phase.
A final remark is that a total derivative term in $\delta(t)$ is usually tolerant in an interferometry experiment because the initial and final position and momentum of different trajectories are the same and they will cancel with each other. It is notable that such a total derivative term is related to the choice of the unitary transform operator $\hat{\mathcal{U}}(t)$. In particular, the operator $\hat{\mathcal{U}}(t)$ allows an arbitrary phase and can be chosen as, for example [62],
$\hat{\mathcal{U}}^{\prime}(t)=\mathrm{e}^{\frac{i}{\hbar}\hat{p}x_{c}(t)}\mathrm{e}^{-\frac{i}{\hbar}\hat{x}p_{c}(t)}=\hat{\mathcal{U}}(t)\mathrm{e}^{\frac{i}{2\hbar}p_{c}(t)x_{c}(t)}.$ (56)
The corresponding pure phase term $\delta(t)$ for $\hat{\mathcal{U}}^{\prime}(t)$ is [63]
$\delta^{\prime}(t)=\frac{1}{2}m\dot{x}_{c}^{2}(t)-\frac{1}{2}m\omega_{0}^{2}x_{c}^{2}(t).$ (57)
Based on the equation of motion (50), one can easily check the equivalence between $\delta(t)$ and $\delta^{\prime}(t)$ up to a total derivative term $\frac{1}{2}m\frac{\mathrm{d}}{\mathrm{d}t}(x_{c}\dot{x}_{c})$, i.e.
$-\frac{1}{2}ma(t)x_{c}(t)=\frac{1}{2}m\dot{x}_{c}^{2}(t)-\frac{1}{2}m\omega_{0}^{2}x_{c}^{2}(t)-\frac{1}{2}m\frac{\mathrm{d}}{\mathrm{d}t}(x_{c}\dot{x}_{c}).$ (58)
Appendix B Witness for Ramsey Interferometry in Spin-1 System
[…]
$\hat{H}_{p}=\hbar\Omega_{p}\begin{pmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{pmatrix}=\hbar\Omega_{p}\hat{T}^{\dagger}\begin{pmatrix}\sqrt{2}&0&0\\ 0&0&0\\ 0&0&-\sqrt{2}\end{pmatrix}\hat{T},$ (59)
where the transform matrix $\hat{T}$ is
$\hat{T}=\begin{pmatrix}\frac{1}{2}&\frac{1}{\sqrt{2}}&\frac{1}{2}\\ \frac{1}{\sqrt{2}}&0&-\frac{1}{\sqrt{2}}\\ \frac{1}{2}&-\frac{1}{\sqrt{2}}&\frac{1}{2}\end{pmatrix}.$ (60)
Then the time-evolution operator $\hat{U}_{p}(t)=\mathrm{e}^{-i\hat{H}_{p}t/\hbar}$ for the pulse acting on the spin-1 system is given by
$\hat{U}_{p}(t)=\hat{T}^{\dagger}\begin{pmatrix}\mathrm{e}^{-i\sqrt{2}\Omega_{p}t}&0&0\\ 0&1&0\\ 0&0&\mathrm{e}^{i\sqrt{2}\Omega_{p}t}\end{pmatrix}\hat{T}=\begin{pmatrix}\cos^{2}\frac{\sqrt{2}\Omega_{p}t}{2}&-\frac{i}{\sqrt{2}}\sin\sqrt{2}\Omega_{p}t&-\sin^{2}\frac{\sqrt{2}\Omega_{p}t}{2}\\ -\frac{i}{\sqrt{2}}\sin\sqrt{2}\Omega_{p}t&\cos\sqrt{2}\Omega_{p}t&-\frac{i}{\sqrt{2}}\sin\sqrt{2}\Omega_{p}t\\ -\sin^{2}\frac{\sqrt{2}\Omega_{p}t}{2}&-\frac{i}{\sqrt{2}}\sin\sqrt{2}\Omega_{p}t&\cos^{2}\frac{\sqrt{2}\Omega_{p}t}{2}\end{pmatrix}.$ (61)
If the time of the microwave pulse is set as $t_{p}=\pi/(2\sqrt{2}\Omega_{p})$, then the corresponding time-evolution operator is
$\hat{U}_{\pi/2}\equiv\hat{U}_{p}\left(t_{p}=\frac{\pi}{2\sqrt{2}\Omega_{p}}\right)=\begin{pmatrix}\frac{1}{2}&-\frac{i}{\sqrt{2}}&-\frac{1}{2}\\ -\frac{i}{\sqrt{2}}&0&-\frac{i}{\sqrt{2}}\\ -\frac{1}{2}&-\frac{i}{\sqrt{2}}&\frac{1}{2}\end{pmatrix}.$ (62)
After the $\pi/2$-pulse, the witness to measure the population of |$\ket{S_{z}=0}$ is simply $\hat{W}^{\prime}=\ket{S_{z}=0}\bra{S_{z}=0}$ |. Since the $\pi/2$-pulse makes a transformation on $\hat{\rho}_{\rm spin}$ to $\hat{U}_{\pi/2}\hat{\rho}_{\rm spin}\hat{U}_{\pi/2}^{\dagger}$, the observable result for the witness $\hat{W}^{\prime}$ is $\mathrm{Tr}\left[\hat{W}^{\prime}\hat{U}_{\pi/2}\hat{\rho}_{\rm spin}\hat{U}_{\pi/2}^{\dagger}\right]$. Therefore, the witness for Ramsey interferometry is given by
$\hat{W}=\hat{U}_{\pi/2}^{\dagger}\hat{W}^{\prime}\hat{U}_{\pi/2}=\begin{pmatrix}\frac{1}{2}&0&\frac{1}{2}\\ 0&0&0\\ \frac{1}{2}&0&\frac{1}{2}\end{pmatrix}.$ (63)
Appendix C Master Equation of Position localisation Decoherence of Single Path
In this appendix, I will derive the master equation for the position localisation decoherence of a single path of the interferometer. The procedure is similar to the derivation of the Redfield master equation which can be found in textbooks such as [64].
Let’s work under the interaction picture where the total Hamiltonian (2) of a single arm can be divided into $\hat{H}_{0}$ and $\hat{H}_{\rm int}(t)=-ma(t)\hat{x}$, where $\hat{H}_{0}$ is of a simple oscillator and $\hat{H}_{\rm int}(t)$ contains all force terms. So the driven force and the gravity signal perform as the mean value of $ma(t)$, while the acceleration noise affects the auto-correlation of $ma(t)$. The dynamical equation for $\hat{\rho}^{I}(t)$ is
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{\rho}^{I}=-\frac{i}{\hbar}[\hat{H}^{I}_{\rm int}(t),\hat{\rho}^{I}(t)].$ (64)
This differential equation can be directly written in an integral form as
$\hat{\rho}^{I}(t)=\hat{\rho}^{I}(0)-\frac{i}{\hbar}\int_{0}^{t}[\hat{H}^{I}_{\rm int}(t^{\prime}),\hat{\rho}^{I}(t^{\prime})]\mathrm{d}t^{\prime}.$ (65)
The basic idea is to do recursion for $\hat{\rho}^{I}(t)$, that is, to insert (65) to the right side of (64), then one may obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{\rho}^{I}=-\frac{1}{\hbar^{2}}\int_{0}^{t}[\hat{H}^{I}_{\rm int}(t),[\hat{H}^{I}_{\rm int}(t^{\prime}),\hat{\rho}^{I}(t^{\prime})]]\mathrm{d}t^{\prime},$ (66)
where $\hat{\rho}^{I}(0)$ is assumed to commute with $\hat{H}^{I}_{\rm int}(t)$, i.e. $[\hat{H}^{I}_{\rm int}(t),\hat{\rho}^{I}(0)]=0$. Besides, the decoherence is usually assumed memoryless, known as the Markovian assumption [25, 64] which means that the $\hat{\rho}^{I}(t)$ can be assumed to only depend on $t$, i.e. the $\hat{\rho}^{I}(t^{\prime})$ in the integral can be substituted by $\hat{\rho}^{I}(t)$. In addition, one can substitute $t^{\prime}$ by $t-\tau$ in the integral, and take the upper limit of the integral to infinite, then obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{\rho}^{I}=-\frac{1}{\hbar^{2}}\int_{0}^{\infty}[\hat{H}^{I}_{\rm int}(t),[\hat{H}^{I}_{\rm int}(t-\tau),\hat{\rho}^{I}(t)]]\mathrm{d}\tau.$ (67)
Further, by substituting the interaction Hamiltonian $\hat{H}^{I}_{\rm int}(t)=-ma(t)\hat{x}^{I}(t)$, and taking the ensemble average, the dynamical equation of the density matrix becomes to
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[\hat{\rho}^{I}(t)]=-\frac{m^{2}}{\hbar^{2}}\int_{0}^{\infty}\mathbb{E}\left[a(t)a(t-\tau)\right]\left[\hat{x}^{I}(t),\left[\hat{x}^{I}(t-\tau),\mathbb{E}[\hat{\rho}^{I}(t)]\right]\right]\mathrm{d}\tau.$ (68)
This is a deformation of the Redfield master equation for a classical stochastic noise. In the original Redfield equation, the environment is assumed as a quantum bath, so the density matrix of the test mass is obtained by tracing out the degrees of freedom of the quantum bath. For the classical random noise, the ensemble average plays a similar role as tracing out, which intimates a correspondence between randomness and quantumness [58, 59].
Note that the time-evolution of the position operator $\hat{x}^{I}(t)$ of a harmonic oscillator satisfies $\hat{x}^{I}(t-\tau)=\hat{x}^{I}(t)\cos\omega_{0}\tau-\hat{p}^{I}(t)/(m\omega_{0})\sin\omega_{0}\tau\sim\hat{x}^{I}(t)\mathrm{e}^{i\omega_{0}\tau}$, so (68) becomes to
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[\hat{\rho}^{I}(t)]=-\frac{m^{2}}{\hbar^{2}}[\hat{x}^{I}(t),[\hat{x}^{I}(t),\mathbb{E}[\hat{\rho}^{I}(t)]]]\int_{0}^{\infty}\mathbb{E}[a(t)a(t-\tau)]\mathrm{e}^{i\omega_{0}\tau}\mathrm{d}\tau.$ (69)
Note that the integral is nothing but $22\pi S_{aa}(\omega_{0})/$ 2 according to the Wiener-Khinchin theorem and the symmetry of $S_{aa}(\omega)$, i.e. $S_{aa}(-\omega)=S_{aa}(\omega)$. So this master equation describes a position-localisation decoherence with a dissipator $\Lambda=\pi m^{2}S_{aa}(\omega_{0})/\hbar^{2}$, and it becomes to the master equation (37) in the main text by switching back to the Schrodinger picture.
Acknowledgements
I would like to thank the China Scholarship Council (CSC) for the financial support. I also thank Guan-Zhuo Yang for the beneficial discussions.